### v1

# How Do Multilingual Models Remember? Investigating Multilingual Factual Recall Mechanisms

Abstract
Large Language Models (LLMs) store and retrieve vast amounts of factual knowledge acquired during pre-training. Prior research has localized and identified mechanisms behind knowledge recall; however, it has primarily focused on English monolingual models. The question of how these processes generalize to other languages and multilingual LLMs remains unexplored. In this paper, we address this gap by conducting a comprehensive analysis of two highly multilingual LLMs. We assess the extent to which previously identified components and mechanisms of factual recall in English apply to a multilingual context. Then, we examine when language plays a role in the recall process, uncovering evidence of language-independent and language-dependent mechanisms.
How Do Multilingual Models Remember?
Investigating Multilingual Factual Recall Mechanisms
Constanza Fierro   Negar Foroutan   Desmond Elliott   Anders Søgaard University of Copenhagen    EPFL
1 Introduction
Large Language Models (LLMs) acquire a vast amount of factual knowledge during pre-training, including many propositional facts of the form “The capital of France is __” Petroni et al. (2019). Such acquisition also occurs in multilingual language models (Kassner et al., 2021; Jiang et al., 2020; Yin et al., 2022), although with considerable variation in performance depending on the input language. This raises an important question: Is this variation caused by the knowledge being stored in a language-specific manner, or is it related to the exact phrasing sensitivity? (Elazar et al., 2021) If trust turns on knowledge recall (Grasswick, 2010; Hawley, 2012; Nguyen, 2022), such assessments will be integral to establishing whether multilingual models are trustworthy.
Mechanistic interpretability aims to understand how and why models process information and make predictions at an internal structural level (Elhage et al., 2022; Nanda et al., 2023; Olah et al., 2020). Recent works have focused on identifying the specific model components responsible for storing acquired knowledge (Meng et al., 2022; Sharma et al., 2024), and understanding how this knowledge is retrieved internally during inference (Geva et al., 2023; Chughtai et al., 2024). However, so far, this research has been exclusively in English and largely limited to autoregressive LMs, except for Sharma et al. (2024), who study these mechanisms in Mamba (Gu and Dao, 2023), a state-space LM.
The acquisition of factual knowledge is usually studied by assessing whether a model predicts the expected fact, given a subject and relation as input, e.g. “[Y], which is the capital of France”. This setup was originally proposed for masked LMs Petroni et al. (2019), where the object to be predicted could be positioned anywhere within the input. With the increasing popularity of decoder-only models (Brown et al., 2020), however, the input format has shifted. For these models, the placeholder (e.g., “[Y]”) has to be placed at the end of the sentence, so that the model predicts the next token, which is the object (Meng et al., 2022; Yu et al., 2023; Geva et al., 2023; Hase et al., 2023; Wang et al., 2024; Cohen et al., 2024). Clearly, this approach presents challenges for languages that do not follow a subject-verb-object (SVO) word order. In languages like Turkish, Japanese, Korean, and Persian, which predominantly use a subject-object-verb (SOV) structure, it can be unnatural and difficult to craft inputs that fit this template for studying acquired factual knowledge.
In this paper, we study the mechanisms of factual recall in multilingual LLMs, focusing on two architectures: the decoder-only model XGLM and the encoder-decoder model mT5; and centering on four key questions: (1) Are the factual recall mechanisms and localization found for English LMs also present in multilingual LLMs? (2) Do these mechanisms function the same way for different languages? (3) Does the model architecture—encoder-decoder or decoder-only—affect these mechanisms? (4) When language is encoded, are there language-dependent mechanisms at play? To address (1)–(4), we first conduct a causal tracing analysis (§3) to determine whether the early MLP modules processing the final subject token are as decisive in multilingual models as in English-centric models (Meng et al., 2022). In XGLM, this effect is absent, while in mT5, we observe a strong causal impact from both the first and last subject tokens across all layers of the encoder.
Second, we analyze the mechanisms described in Geva et al. (2023), with a focus on when information from the subject and relation flows to the last token, and when the object is extracted from the subject representation (§4). We find that information flows similarly in monolingual and multilingual decoder-only models, while for mT5, there are discrepancies. In terms of the extraction event, in XGLM, both feed-forward and attention sublayers contribute to it, in contrast to English autoregressive models, where attention modules do most work. In mT5, this mechanism is primarily performed by the cross-attention, occurring later in the process and in a more localized manner compared to decoder-only models.
Finally, we use activation patching (Zhang and Nanda, 2024; Ghandeharioun et al., 2024) and insert the last token representation from an English input into the forward pass of another language (§5); this way, we are able to characterize more fully the multilingual recalling process. We find that in XGLM there are two distinct phases. First the relation representation propagates to the last token, and then the output language. While for mT5, both the relation and language propagate in the later layers, with the relation being encoded slightly earlier. Additionally, we show that the subject and relation representations are language-independent, as they can be interchanged and used in different languages; while the extraction event is language-specific. The findings of this work contribute to the understanding of crosslingual factual knowledge transfer and consistency, and can further inform knowledge editing methodologies, and trustworthiness assessments in multilingual models.
2 Experimental Setup
We focus on the basic form of factual knowledge recall, where LMs are tasked with predicting the correct object $o$ for a given subject $s$ and a relation $r$. These $(s,r,o)$ triplets are obtained from WikiData, and natural language templates are used to describe the relation, with placeholders for the subject and object. For the relation born-in, a template could, e.g., be “[X] was born in [Y]”, where [X] represents the placeholder for the subject and [Y] is the object to be predicted by the model.
We use the mParaRel dataset (Fierro and Søgaard, 2022), with triplets and templates for 45 languages. [^1] We augment mParaRel by fetching WikiData aliases of the objects to be able to identify more cases of correct object prediction. We filter out trivial examples where the object is directly mentioned in the query, e.g., South Asia belongs to the continent of. For our analysis, we select 10 typologically diverse languages representing a variety of scripts, language families, and word orders, including SVO, SOV, and VSO structures: English (en), Spanish (es), Vietnamese (vi), Turkish (tr), Russian (ru), Ukrainian (uk), Japanese (ja), Korean (ko), Hebrew (he), Persian (fa), and Arabic (ar).
We analyze two different architectures of multilingual language models, XGLM (Lin et al., 2021),
Language XGLM mT5
en 1812 1543
es 1380 1192
vi 1646 993
tr 418 1058
ru 830 683
uk 213 456
ko 116 630
ja 6 358
he 13 565
fa 7 406
ar 811 488
Table 1: Number of memorized knowledge triplets $(s,r,o)$.
a decoder-only architecture with 7.5B parameters and 32 layers, and mT5-xl (Xue et al., 2021), an encoder-decoder architecture of 3.7B parameters and 24/24 encoder/decoder layers. For XGLM, we only use the templates in mParaRel that have the object placeholder at the end of the sentence, while for mT5, we use all the templates available. XGLM pre-training data includes 30 languages but it does not include uk, he, and fa; while mT5 includes 101 languages and covers all the ones studied here.
As we aim to investigate the process of knowledge recall, and similarly to previous work (Meng et al., 2022; Geva et al., 2023), we only consider examples where the model predicts the correct object completion. Since mParaRel provides multiple templates (paraphrases) for each relation, we greedy-decode for every available template corresponding to a given triplet and check for an exact match. [^2] We perform greedy decoding of a maximum of 50 tokens and check for exact match of any of the object aliases. In practice, the 50 tokens generation is an overestimate, as on average, around 0.5 tokens are generated prior to the target object (for some templates an article or other filling tokens are generated before predicting the object). If multiple templates trigger the correct object, we randomly select one for the analyses. Once an exact match is found, any extra tokens generated before the object are added to the example input text. This ensures that when we feed again the example to the model, the next token predicted is the first token of the object. [^3] For mT5, we feed the mParaRel input into the encoder with a sentinel token in the object placeholder. In the decoder, we provide the beginning-of-sequence token followed by the sentinel token. We only check the tokens generated next to the first sentinel token, as the pre-training task of the model is to generate the text for each sentinel in the input, the decoder usually continues to generate answers for other sentinel tokens. Any tokens generated before the target object, are added to the decoder tokens since adding these to the encoder does not ensure that the next token predicted will be the object. Table 1 presents the number of examples for which the correct object is predicted. Due to the limited number of examples memorized by XGLM in ja, he, and fa, we exclude these languages from our main analysis of XGLM.
3 Causal Tracing
We first analyze which hidden states in the model’s computation are more important than others when recalling a fact. Following Meng et al. (2022), we trace the causal effects of hidden states using causal mediation analysis (Pearl, 2022). To do this, we feed the model a corrupted input and selectively restore a particular component in the computational graph, such as the attention output of layer $l$. We then measure how much of the original probability (from the uncorrupted input) is recovered, thereby assessing the indirect influence of the restored component on the model’s prediction. Specifically, we corrupt the input by adding Gaussian noise to the subject tokens, [^4] We follow Meng et al. (2022) and add $\epsilon\sim\mathcal{N}(0,(3\sigma)^{2})$, with $\sigma$ being the standard deviation of the subjects tokens embeddings from the data used. and observe the probability $\hat{\mathbb{P}}$ of the originally predicted token. Then, we run inference again on the corrupted input, but this time, we restore a specific hidden state in the model and track the probability ^$\hat{\mathbb{P}}_{\text{restored}}$ of the originally predicted token. We study the indirect effect $\text{IE}=\hat{\mathbb{P}}_{\text{restored}}-\hat{\mathbb{P}}$.
Consider a transformer model with $L$ layers; let $h_{t}^{l}$ be the representation of the token $t$ at layer $l$. When the model is an encoder-decoder, let $e_{i}^{l}$ be the representation of the encoder layer $l$ for the $i$-th token in the input. Then, the encoder layer computes $h_{t}^{l+1}=h_{t}^{l}+s^{l}+f^{l}$ and the decoder $h_{t}^{l+1}=h_{t}^{l}+s^{l}+c^{l}+f^{l}$, where $s^{l}=\text{Self Attn.}(h_{0}^{l}\ldots h_{t}^{l})$, $c^{l}=\text{Cross Attn.}(h_{t}^{l},e_{0}^{l}\ldots e_{n}^{l})$ and $f^{l}=\text{MLP}(h_{t}^{l}+s^{l}+c^{l})$. If decoder-only, then $c^{l}$ does not apply.
Let $h_{t}^{l}$ be the hidden state from the clean run and $\overline{h}_{t}^{l}$ that of the corrupted run. We restore a state by setting $\overline{h}_{t}^{l}\leftarrow h_{t}^{l}$, and similarly, we restore the self-attention layers contribution by setting $\overline{s}^{l}\leftarrow s^{l}$ for all the attention modules in a window of size $w$ (analogous for $c$ and $f$). [^5] We use windows because, generally, the contributions of the sub-layers are gradual (Geva et al., 2021). In other words, multiple layers contribute the same behavior to the residual, and their sum produces the observed effect. Following Meng et al. (2022), we set $w=10$ for XGLM, and $w=6$ for mT5 (as it has 24 layers). We repeat the corrupted run with restoration ten times, each with different noise samples, and report the average.
We compare our results in multilingual LLMs to those of Meng et al. (2022), who analyzed the factual recall of GPT-2 XL in English, [^6] They used the CounterFact dataset, which is based on the ParaRel dataset. Thus, the English data is the same as the one we use here. and reached two key conclusions. First, they identified an “early site”, where the MLPs processing the last subject token in the early and middle layers play a crucial role in recovering from input corruption. Second, they found a “late site”, where the attention modules processing the last token in the later layers also significantly contribute to prediction recovery. [^7] While they consider the late site unsurprising, as it directly precedes the final prediction, this observation primarily applies to hidden state restoration, not necessarily attention restoration. We, however, interpret the late-site attention as aligning with what Geva et al. (2023) referred to as the extraction event (§4).
We present the causal analysis plots for each model and language in Appendix B and present our main observations here. We do not find for XGLM an early site of MLP modules nor for English or any of the other languages. Thus, we conclude that this feature is specific to English-centric models, as Sharma et al. (2024) also observed it for another English-centric architecture (Mamba). We do observe a causal effect of the last subject token when restoring the state in the early layers. This is somewhat expected as it reverts the corruption performed in the input, but it is important to note until which layer it is possible to recover from this corruption only by restoring the last subject token. It also highlights the relevance of the last subject token in decoder-only models for all languages.
Figure 1: Attention knockout between the last token and a given set of tokens. Each layer represents the effect of the knockout on a window of $w$ layers. Top XGLM ($w=6$), bottom mT5 ($w=4$).
For mT5, we do observe a causal effect from MLP modules of the subject tokens, but it is not an early site as in GPT. Rather it is a causal effect along all the layers of the encoder, in both the first and last token of the subject, with some languages having a stronger effect in the last subject token. In terms of late site, we find it in the self attention for XGLM, and in the cross attention layers in mT5. For both mT5 and XGLM, we also observe a late site of the MLP at the last token, though this effect is smaller than that of the attention or the subject.
We conclude that the multilingual models behave in a similar way for the different languages, finding that the causal effects of both MLP and attention modules are positioned the same, and that for decoder-models the last subject token (at the early layers) is critical for the recalling, while for encoder-decoder models the MLPs processing both the first and last subject token encode the necessary information (at any encoder layer).
4 Factual Recall Components
Geva et al. (2023) described the process of factual knowledge recall in English autoregressive models as a three-step mechanism: (a) the subject representation is enriched (i.e., related attributes are encoded); (b) the relation and subject information are propagated to the last token; and (c) the final predicted attribute is extracted by attention layers. We analyze the information flow to the last token, and then the extraction of the predicted attribute.
Information Flow
Like Geva et al. (2023), we use attention knockout to visualize the propagation of information to the last token. Based on insights from the causal analysis, we intervene in the last token self-attention in XGLM and cross-attention for mT5. Specifically, we set the attention score to be zero between the last token and another set of tokens $\{t\}$. Given that the contributions of attention blocks are incremental, we perform this intervention within a window of size $w$ around layer $l$. The set of tokens $\{t\}$ consists of (1) subject tokens, (2) non-subject tokens (relation tokens), and (3) the last token (in mT5, the last token refers to the last token in the decoder input). We use a window of size 6 for XGLM and of size 4 for mT5, following Geva et al. (2023) to knock out around 18% of layers. We then measure the change in probability of the originally predicted first token (without knockout), aggregating the results across examples by calculating the relative difference (i.e., the change in probability proportional to the original probability). Thus, a big drop in relative probability indicates a critical flow of information at that layer, from the selected tokens to the last token.
We present the results in Figure 1 (plots per language in Appendix C). For XGLM model, we find broadly similar dynamics to those of GPT (Geva et al., 2023) with a critical flow from the subject at later layers around 25, and negligible flow from the last token to itself. However, there is a strong propagation of information from the non-subject tokens throughout the early and middle layers, shown by the drop of -80% in probability, which reaches -100% for some languages in layers 5-10.888GPT2-xl attention to non-subject tokens is more pronounced in the early layers (-40%), decreasing from there. GPT-J attends to non-subject tokens through early and mid layers but with a weaker drop (-20%) (Geva et al., 2023). Most languages follow a similar pattern, but ru benefits from the knockout to the subject token in early layers; for ar and ko, XGLM attends strongly to non-subject tokens until the end. The mT5 results show a different behavior. First, the last token does attend to itself, a pattern that is more pronounced in the early and later layers, but less so around the middle layers. [^9] We hypothesize that this is probably due to the last token encoding information about which sentinel token is generated at this step (indicating where in the input it should fill), since we see an almost identical curve for all the languages when the last token cannot attend to itself or when it cannot attend to the sentinel token in the decoder (Figure 14). Second, the last token attends to non-subject tokens across all layers, with a weaker drop at the early layers compared to XGLM. Finally, similar to XGLM, there is critical information flowing from the subject token in the later layers. However, this drop is larger and occurs even later in the computation compared to XGLM. Most languages behave similarly, with the exception of ja and ko, exhibiting less flow from last and non-subject tokens in the early layers.
Figure 2: Extraction rates. Left XGLM, right mT5.
Prediction Extraction
Like Geva et al. (2023), we measure extraction events at each layer of the model. Using the notation from §3, let $h^{l}$ be the representation of the last token at layer $l$, and let $E$ be the embedding matrix, then the predicted token is $o^{*}=\arg\,\max(Eh^{L})$. We say there is an extraction event at the self-attention at layer $l$ if $\arg\,\max(Es^{l})=o^{*}$ (similarly for $c^{l}$ and $f^{l}$). The extraction rate is the proportion of examples for which an extraction event occurs at a given layer.
Figure 2 shows the extraction rates for both models, with a detailed language breakdown in Appendix D. A key difference between XGLM and GPT-2 (Geva et al., 2023) is the prominence of MLP modules in the extraction of the object in XGLM. We note that this is not simply because the MLP is forwarding the already extracted object from the preceding attention layer. We plot in Figure 17 how often an MLP has an extraction event while the preceding attention has not, and we find a similar plot to the one in Figure 2, where the MLP has a higher peak in layer 25 than the attention. Additionally, we observe that for English (Figure 15), in line with the results for GPT-2, the attentions have higher extraction rate than the MLPs.
For mT5, the cross-attention layers are the main mechanism for extraction, although the MLP layers also contribute to a lesser degree, resembling GPT-2’s behavior. However, for certain languages (he, fa, ko, and ru), the MLPs show comparable or even higher extraction rates (Figure 16). Finally, we observe that in mT5, extraction events occur later in the computation, with peaks at layers 19 and 22.
5 Information Flow Through Patching
Finally, we study the role of language in the recall prediction of transformer models. We employ activation patching (Wang et al., 2022) by intervening in the model’s computation to identify critical steps. Through these analyses, we contribute three key elements to the understanding of the recall process in transformer models: (1) we localize exactly at which layer the relation information flows to the last token; (2) we identify the extraction mechanism in multilingual LLMs; and (3) we explore how language plays a role in recall prediction.
Figure 3: Patching strategy. The patch example (left) and the context example (right). The patch of this transformer is performed at layer 0 and it affects the computation of layer 1, thus the output is perturbed.
Let $h^{l}$ be the representation at layer $l$ of the last token in the input. For a given input $p$, we take $h^{l}_{p}$ and patch it into the forward pass of another input $c$ at the same layer; that is, we set $h^{l}_{t,c}\leftarrow h^{l}_{t,p}$, and then continue the forward pass. We refer to the example $p$ as the patch example, and $c$ as the context example (See Figure 3). By patching the last token representation in each layer and measuring the model’s predicted output, we can study when the representation contains the information about relation, language, and predicted object, and how cross-lingual these representations are.
An input example expresses a relation $r$ and a subject $s$ in language $\mathcal{L}$. We conduct three experiments where the patch and context examples share certain input characteristics. In particular, we test three different conditions: (1) $\{=\mathcal{L},\neq r,\neq s\}$, where the language is the same but the relation and subject differ; (2) $\{\neq\mathcal{L},=r,\neq s\}$, where the relation remains the same but the language and subject differ; and (3) $\{\neq\mathcal{L},\neq r,=s\}$, where the subject is the same but the language and relation differ. (1) allows us to study how the relation is encoded while controlling for the language, whereas (2)-(3) help us explore how the interaction between relation and language affects the extraction of the object from the subject. In all experiments, English is used as the patch language, $\mathcal{L}_{p}=\text{en}$.
Let the input of the patch be $(\mathcal{L}_{p}(r_{p}),\mathcal{L}_{p}(s_{p}))$ and the input of the context be $(\mathcal{L}_{c}(r_{c}),\mathcal{L}_{c}(s_{c}))$. When there is no intervention, the model correctly predicts $\mathcal{L}_{p}(o_{r_{p},s_{p}})$ and $\mathcal{L}_{c}(o_{r_{c},s_{c}})$ respectively. [^10] For simplicity, we may refer to the object as $\mathcal{L}_{p}(o_{p})$ or $\mathcal{L}_{c}(o_{c})$ when $r$ and $s$ are from the same input. To analyze the effect of patching, we study both the output probabilities and the final predictions. First, we evaluate the probability of the originally predicted token, $\mathcal{L}_{p}(o_{p})$ and $\mathcal{L}_{c}(o_{c})$, and similar to §4 we aggregate across examples by calculating the relative difference (the change in probability proportional to the original (unpatched) probability). Second, we examine the predicted token, and we check whether it corresponds to $\mathcal{L}_{p}(o_{c})$, the context answer but in the patch language, or to $\mathcal{L}_{c}(o_{p})$, the patch answer but in the context language. Note that we consider only pairs of patch-context examples for which we have the corresponding objects, and which do not share the same spelling. [^11] We limit ourselves to pairs where: (1) mParaRel has the expected object in the other language (e.g. that we have the aliases of the context object in the patch language $\mathcal{L}_{p}(o_{c})$); and (2) the prediction can only match one of the object aliases sets, e.g. to say that the predicted token $t=\mathcal{L}_{p}(o_{c})$ then $t\neq\mathcal{L}_{c}(o_{c})$, which may happen if the languages share the spelling of the object $o_{c}$. For example, “Asia” is spelled the same in English and Spanish. As a result, the number of examples varies when we study output probabilities, predictions of $\mathcal{L}_{c}(o_{p})$, or predictions of $\mathcal{L}_{p}(o_{c})$. See Table 2 for the number of examples for each model and experiment.
Different Relation, Different Subject
First, we analyze the pairs of examples with $\{=\mathcal{L},\neq r,\neq s\}$. [^12] For example, $r_{p},s_{p}=$“The capital of France is” and $r_{c},s_{c}=$“The language spoken in Germany is”. Then, $o_{r_{p},s_{c}}=$ Berlin and $o_{r_{c},s_{p}}=$French. We study $\sim$1000 examples for which mParaRel has the objects $o_{r_{c},s_{p}}$ and $o_{r_{p},s_{c}}$. [^13] The total number of examples that have different spellings when checking for $o_{r_{p},s_{c}}$ is 984 (994) in XGLM (mT5), and for $o_{r_{c},s_{p}}$ is 974 (990). We present the prediction results in Figure 4, with corresponding probability plots in Figure 20. We observe that the relation information is propagated to the last token representation by layer 13 in XGLM and layer 16 in mT5, since the patching at that point starts producing the object corresponding to the relation in the patch $o_{r_{p},s_{c}}$. By layer 22 in XGLM and layer 18 in mT5, we observe that the object has been fully extracted and is encoded within the last token representation, as predictions from these layers correspond to $o_{r_{p},s_{p}}$. Layer 22 in XGLM aligns with the first attention peak in the extraction rate of English (Figure 15), and layer 18 in mT5 matches the first cross-attention peak (Figure 16). This shows that the extraction measured by the vocabulary projection indeed represents the point at which the object is extracted and encoded in the last token representation. If the object was encoded earlier and only decoded with the extraction event, we would see $o_{r_{p},s_{p}}$ being predicted from patches at earlier layers.
Figure 4: Counts of examples where the patch produces the object prediction. Top is XGLM, bottom mT5.
Same Relation, Different Subject
We have localized the layers where the relation representation flows to the last token. Now, we analyze when the language of the input text propagates to the last token. Is the language information entangled to the subject or the relation representation? Or is the language encoded separately from these two? Here, we study pairs of examples with $\{\neq\mathcal{L},=r,\neq s\}$. [^14] For example: “France’s capital is” and “La capital de Alemania es” (Gloss: “The capital of Germany is”). If we obtain $\mathcal{L}_{p}(o_{c})$, this suggests that the language is encoded in the last token representation before the extraction happens. On the other hand, if we obtain $\mathcal{L}_{c}(o_{p})$, we could infer that the language is encoded after the extraction.151515 After filtering examples with similar object spellings, the total number of pairs we have per language is around 150-800.
We present the probability of the context answer token $\mathcal{L}_{c}(o_{c})$ in Figure 5 (the probability of $\mathcal{L}_{p}(o_{p})$ and per language plots are provided in Appendix E.2). First, we observe that the probability of $\mathcal{L}_{p}(o_{p})$ increases sharply around layers 21-25 for XGLM and around 18-21 for mT5. For XGLM, regarding the probability of $\mathcal{L}_{c}(o_{c})$, we observe that patching in the early layers generally hurts the model’s performance for most languages. However, patching in the middle layers (10-20) either increases the probability or results in only a minimal decrease. Given that the relation is the same in both examples, this suggests that by these middle layers, $r$ is encoded in $h^{l}_{p}$, and in the subsequent layers, the subject and language of the context are integrated to yield the final prediction, $\mathcal{L}_{p}(r)+\mathcal{L}_{c}+s_{c}=\mathcal{L}_{c}(o_{c})$. Moreover, for some languages the relation representation from the patch is better than the one constructed using the context input, as the relative probability is positive.
Figure 5: Relative probability change of $\mathcal{L}_{c}(o_{c})$ when patching the last token representation between pairs of examples with the same relation and a different subject (Top XGLM, bottom mT5).
In the case of mT5, however, the probability of $\mathcal{L}_{c}(o_{c})$ consistently decreases, plateauing in the middle layers before dropping to zero by layer 22. From the attention knockout analysis, we recall that in mT5, the subject is integrated into the last token representation only after layer 15, while the relation is consistently represented throughout the middle layers. Therefore, these patching results imply that the relation encoded in the last token from the patch input does not help in retrieving the correct context object in the target language in mT5, whereas it proves useful in XGLM.
Figure 6: Histogram of patches that cause the context answer to be predicted in the patch language, $\mathcal{L}_{p}(o_{c})$ (Top XGLM, bottom mT5).
In terms of predicted tokens, we find that across-the-board $\mathcal{L}_{c}(o_{p})$ is rarely predicted while $\mathcal{L}_{p}(o_{c})$ appears frequently (Table 3). This suggests that the last token representation encodes the patch language $\mathcal{L}_{p}$ but not yet the object, as the object is derived from the subject context. Figure 6 shows the layers where $\mathcal{L}_{p}(o_{c})$ is predicted (plot for $\mathcal{L}_{p}(o_{c})$ in Figure 31). As discussed in §4, the extraction event typically happens around layers 18-28 in XGLM and 16-22 in mT5. Therefore, we observe that the language is attended to at the beginning of the extraction phase, because if we patch earlier, the output is in the context language. Then, the language is not attended anymore but rather the language information that already flowed to the last token is the one used for the rest of the extraction phase. We can interpret the last token representation as containing the “task” (Todd et al., 2024), the relation that needs to be extracted and in which language, which is used in the extraction event. This task vector can be transferred to contexts in another language in both models, as the task vector is constructed from the patch input and used in the context subject representation to predict $\mathcal{L}_{p}(o_{c})$.
Different Relation, Same Subject
We just saw that the representation from the patch will encode at some point the output language but not yet the object. In the previous experiment both examples had the same relation, so it could be that we observe the prediction $\mathcal{L}_{p}(o_{r,s_{c}})$ because the relation is language specific and encodes the output language. To analyze if this is the case, we now apply patching on examples with different languages and relations but the same subject $\{\neq\mathcal{L},\neq r,=s\}$. [^16] For example: “[Y]’s capital is” and “El idioma oficial de [Y] es” (Gloss: “The official language of [Y] is”). The number of examples is around 150-300 per language, while for XGLM in uk and tr is around 30-70 and only 6 in ko (Table 2). When analyzing the predictions of $\mathcal{L}_{c}(o_{p})$ and $\mathcal{L}_{p}(o_{c})$, as we restrict the examples to have different spellings, the total number of pairs is around 75-250 depending on the language.
Figure 7: Relative probability change of $\mathcal{L}_{c}(o_{c})$ when patching the last token representation between pairs of examples with the different relation and the same subject (Top XGLM, bottom mT5).
We present the probability of the original token $\mathcal{L}_{c}(o_{c})$ and $\mathcal{L}_{p}(o_{p})$ in Figures  7 and 26, respectively. We observe that the probability of $\mathcal{L}_{p}(o_{p})$ starts increasing in layers 10-15 for es, vi, tr, and ko while the increase starts from layers 15-20 for ru, uk, and ar. In mT5, this increase is more uniform across languages and occurs slightly later. As for $\mathcal{L}_{c}(o_{c})$, unlike the previous experiment, we observe that in XGLM, the probability decreases early for all languages (except tr), plateaus around the middle layers, and then drops to zero by the mid-layer range. In mT5, similar to the previous experiment, the probability of $\mathcal{L}_{c}(o_{c})$ drops at the beginning but, instead of plateauing, it continues to decline until it reaches zero around layer 17.
In terms of predictions, we find that $\mathcal{L}_{c}(o_{p})$ is frequently predicted for XGLM (Figure 8), while $\mathcal{L}_{p}(o_{c})$ appears but less often (see Figure 31 and Table 3). This suggests that in layers 15-21, the patched last token representation primarily captures the relation $r_{p}$, without yet encoding the output language or subject information. Therefore, in XGLM, the relation and language representations are disentangled, as the relation flows to the last token before the output language does. Allowing the relation to be combined with different languages.
As for mT5, we observe very few examples where $\mathcal{L}_{c}(o_{p})$ or $\mathcal{L}_{p}(o_{c})$ are predicted, which aligns with the findings of the two former experiments, where we see that the relation is encoded in the last token in layers 15-21 (Figure 4), and the language flows to the last token around layer 15-19 (Figure 6). We conclude that both the relation and language flow to the last token around the same time, and thus, in this experiment, we cannot see a disentangled behavior. This presents an interesting contrast with XGLM. The decoder in mT5 has access to the same encoder representations throughout all its layers, so it does not need (and thus does not learn) to attend to these earlier or in different stages. By contrast, in a decoder-only model like XGLM, the last token has access to representations that evolve across layers, so it learns to attend to relevant information when it becomes most salient. Nonetheless, we cannot reach a definite conclusion on whether the language representation in mT5 is entangled or not to the relation representation. More detailed analysis of what is being attended when the relation flows and when the language flows should be performed in future work.
Figure 8: Histogram of patches that cause the prediction to be $\mathcal{L}_{c}(o_{p})$.
6 Conclusion
In this paper, we analyzed the mechanisms involved in the recall of factual knowledge in 10 languages using multilingual transformer-based LMs. We compare these findings to previous research on English recall in English autoregressive LLMs. We discovered that some mechanisms, such as the flow of subject representations in the later layers and the extraction phase, are shared between multilingual and monolingual models. However, we also identified notable differences, including the absence of knowledge localization in the early MLPs and the joint role of MLPs and attention modules during the extraction phase in multilingual models. Finally, we also studied when language encoding occurs during the recall process by patching representations. In XGLM, we observed two distinct phases: initially, the relation flows to the last token, followed by the language. In contrast, in mT5, both the relation and language flow to the last token at similar layers. This suggests that while the relation and subject representations are multilingual and can be used to extract objects across languages, the extraction phase itself appears to be language-specific, as the language is encoded prior to extraction. These findings provide new evidence to understand the factual knowledge recall in transformer LMs, contributing with analyzes in different architectures and languages, and suggesting new directions for the study of cross-lingual transfer.
7 Limitations
In this paper, we examined two model architectures, leaving out the effects of model sizes, instruction fine-tuning, or models like Llama that can behave multilingually but have less coverage and less multilingual pre-training data. Additionally, our analyses were conducted on 500-1000 examples per language, which we believe provides a sufficient sample size for generalization; however, the results are inherently limited by the relations present in the mParaRel dataset, which may not capture all factual nuances. Additionally, although we analyzed 10 diverse languages, many more languages exist, and further research is needed to confirm the generalizability of our findings across a broader linguistic spectrum. Lastly, we described the main mechanisms found in XGLM and mT5, however other weaker mechanisms could be at play, which could describe, for example, the low extraction rates found for some languages (Figure 16) or the few examples where the object seems to be encoded from early layers before the extraction takes place (Figure 31).
Appendix A Related Work
Factual Knowledge Recall.
Petroni et al. (2019) analyzed the factual knowledge encoded in pre-trained language models, leading to further multilingual research. [^17] One may object to calling the ability to solve cloze tests as knowledge. LLMs are sometimes inconsistent and cannot always justify the propositions they generate. We side with Fierro et al. (2024) in thinking knowledge may nevertheless be the appropriate concept, since LLMs are mostly consistent and can sometimes provide justification, e.g., in virtue of world models or training data attribution. Kassner et al. (2021); Jiang et al. (2020) extended this work by translating the LAMA dataset (Petroni et al., 2019) and evaluating the factual knowledge acquired by multilingual pre-trained models, while Wang et al. (2024) evaluated it from the perspective of knowledge edits (Zhu et al., 2021). Building on this, Yin et al. (2022) curated a dataset that focuses on the cultural aspect of factual knowledge. While Fierro and Søgaard (2022) presented the mParaRel dataset181818mParaRel is a machine-translated version of the English templates from ParaRel (Elazar et al., 2021), where subjects and objects are translated using WikiData identifiers. to examine the consistency of factual predictions in multilingual pre-trained language models. In a related effort, Qi et al. (2023) measured the cross-lingual consistency of multilingual LLMs by evaluating the consistency across languages, rather than just across paraphrases in the same language.
Interpretability on Factual Knowledge Recall.
As discussed earlier, Meng et al. (2022) was among the first to analyze, from a mechanistic perspective, whether a model’s knowledge could be localized; while Geva et al. (2023) conducted the first study on how information flows within a model to construct the final predicted object representation. Both studies focused on the GPT models (Brown et al., 2020), using English-language data. Later, Chughtai et al. (2024) contributed by demonstrating that models arrive at an answer through several independent components, with the sum of these components yielding the correct object. Their analysis covered both GPT and Pythia models, also using English data. Subsequently, Sharma et al. (2024) applied the mechanisms identified by Meng et al. (2022) and Geva et al. (2023) to the state-space language model Mamba (Gu and Dao, 2023), focusing on facts expressed in English. Additionally, Hernandez et al. (2024) examined whether a linear transformation of the subject is sufficient to represent various relations, analyzing GPT and Llama models with English data.
Activation Patching.
Ghandeharioun et al. (2024) introduced the Patchscope framework to study what an intermediate representation encodes. In their experiments, they focused on “decoding” the contents of a latent representation by patching it into the forward pass of another input. Similarly, Zhang and Nanda (2024) explored the differences and implementation nuances of patching representations versus adding Gaussian noise. More relevant to the type of patching we conducted in this work, Dumas et al. (2024) investigated a word translation task to assess whether the model encodes concepts that are disentangled from the language during translation. This builds on earlier findings by Wendler et al. (2024), who observed that models like Llama-2 (Touvron et al., 2023) seem to use English as an internal pivot language. Consistent with our results, Dumas et al. (2024); Foroutan et al. (2022) provided evidence for the existence of language-agnostic representations.
Multilingual Language Models
In this work, we focused on generative models and selected mT5 (Xue et al., 2021) and XGLM Lin et al. (2021) due to their broad language coverage, highly multilingual training data, and pre-training with up-sampling of low-resource languages to ensure more balanced language representation. However, many other models are available. For instance, BLOOM (Le Scao et al., 2023) is a decoder-only model trained on 46 languages, although low factual recall accuracy has been reported (Qi et al., 2023). Llama-2 (Touvron et al., 2023) has been used in some multilingual analyses (Wendler et al., 2024; Dumas et al., 2024), but it is primarily an English-centric model given that 90% of its pre-training data is in English. More recent models that could have been considered include mGPT (Shliazhko et al., 2022) and EuroLLM (Martins et al., 2024), both decoder-only models trained on 61 and 34 languages, respectively. On the other hand, the Aya-101 model (Üstün et al., 2024) is an instruction fine-tuned version of mT5; however, we focus on pre-trained multilingual LMs to compare our analysis to the English monolingual pre-trained versions previously studied.
Appendix B Causal Tracing
Figure 9: XGLM causal analysis for each language (rest of the languages in Figure 10). Average probability (of the originally predicted token) recovered when corrupting the subject in the input and restoring the hidden representation at a given layer (State), or restoring the MLP / Self Attn. in a window of 10 layers.
Figure 10: XGLM causal analysis for each language (Rest of the languages in Figure 9). Average probability (of the originally predicted token) recovered when corrupting the subject in the input and restoring the hidden representation at a given layer (State), or restoring the MLP / Self Attn. in a window of 10 layers.
Figure 11: Causal analysis of mT5 for each language (Rest of the languages in Figure 12). Average probability (of the originally predicted token) recovered when corrupting the subject in the input and restoring the hidden representation at a given layer (State), or restoring the MLP / Self Attn. / Cross Attn. in a window of 6 layers.
Figure 12: Causal analysis of mT5 for each language (Rest of the languages in Figure 11). Average probability (of the originally predicted token) recovered when corrupting the subject in the input and restoring the hidden representation at a given layer (State), or restoring the MLP / Self Attn. / Cross Attn. in a window of 6 layers.
Appendix C Attention Knockout
Figure 13: XGLM attention knockout for each language.
Figure 14: Attention knockout for each language in mT5. We present the results from knocking out the information from the last token in the decoder (last) and the masked object in the encoder (enc_sentinel). However, from the encoder sentinel there is no much flow of information so these are not plot in the main body of the paper.
Appendix D Extraction Event
Figure 15: XGLM extraction rates for each language.
Figure 16: mT5 extraction rates for each language.
Figure 17: Number of extraction events split by precedence (or not) of an extraction event in the self-attn (cross-attn). Left XGLM, right mT5.
Figure 18: Number of extraction events split by precedence (or not) of an extraction event in the self-attn in XGLM for each language.
Figure 19: Number of extraction events split by precedence (or not) of an extraction event in the cross-attn in mT5 for each language.
Appendix E Patching
XGLM mT5
$=$ Relation, $\neq$ Subject $\neq$ Relation, = Subject $=$ Relation, $\neq$ Subject $\neq$ Relation, = Subject
Source - Target All $o_{s}^{l_{s}}\neq o_{s}^{l_{t}}$ $o_{t}^{l_{s}}\neq o_{t}^{l_{t}}$ All $o_{s}^{l_{s}}\neq o_{s}^{l_{t}}$ $o_{t}^{l_{s}}\neq o_{t}^{l_{t}}$ All $o_{s}^{l_{s}}\neq o_{s}^{l_{t}}$ $o_{t}^{l_{s}}\neq o_{t}^{l_{t}}$ All $o_{s}^{l_{s}}\neq o_{s}^{l_{t}}$ $o_{t}^{l_{s}}\neq o_{t}^{l_{t}}$
en-es 849 155 140 325 75 82 842 830 779 263 250 249
en-vi 862 369 377 337 177 180 854 844 823 209 207 204
en-tr 974 260 253 68 36 15 815 802 794 209 206 205
en-ru 823 565 588 218 187 177 796 795 787 191 191 190
en-uk 914 734 715 35 32 32 853 850 852 136 136 133
en-ko 995 993 991 6 6 6 744 744 743 165 165 165
en-he 962 903 458 0 0 0 778 778 778 166 166 166
en-fa 537 528 513 0 0 0 714 714 712 114 114 114
en-ar 829 818 816 212 209 211 775 775 774 159 159 159
Table 2: Total number of pairs of examples considered in the patching experiments.
XGLM mT5
$=$ Relation, $\neq$ Subject $\neq$ Relation, $=$ Subject $=$ Relation, $\neq$ Subject $\neq$ Relation, $=$ Subject
Patch - Context $\mathcal{L}_{c}(o_{r,s_{p}})$ $\mathcal{L}_{p}(o_{r,s_{c}})$ $\mathcal{L}_{c}(o_{r_{p},s})$ $\mathcal{L}_{p}(o_{r_{c},s})$ $\mathcal{L}_{c}(o_{r,s_{p}})$ $\mathcal{L}_{p}(o_{r,s_{c}})$ $\mathcal{L}_{c}(o_{r_{p},s})$ $\mathcal{L}_{p}(o_{r_{c},s})$
en-es 1.3% (2) 38.6% (54) 49.3% (37) 7.3% (6) 0.7% (6) 21.6% (168) 0.8% (2) 10.4% (26)
en-vi 0.5% (2) 67.1% (253) 11.3% (20) 13.3% (24) 0.2% (2) 20.4% (168) 1.0% (2) 13.7% (28)
en-tr 1.2% (3) 52.6% (133) 47.2% (17) 33.3% (5) 0.2% (2) 26.3% (209) 2.4% (5) 11.2% (23)
en-ru 2.3% (13) 53.1% (312) 63.6% (119) 6.2% (11) 1.4% (11) 36.1% (284) 7.9% (15) 11.6% (22)
en-uk 1.1% (8) 26.2% (187) 15.6% (5) 18.8% (6) 0.4% (3) 31.8% (271) 4.4% (6) 9.0% (12)
en-ko 0.4% (4) 45.0% (446) 83.3% (5) 33.3% (2) 0.5% (4) 32.3% (240) 4.2% (7) 13.3% (22)
en-he 0.0% (0) 57.9% (265) - - 0.3% (2) 38.4% (299) 5.4% (9) 16.3% (27)
en-fa 0.0% (0) 42.3% (217) - - 0.8% (6) 25.3% (180) 3.5% (4) 16.7% (19)
en-ar 8.8% (72) 39.7% (324) 40.7% (85) 4.7% (10) 0.9% (7) 32.0% (248) 2.5% (4) 9.4% (15)
Table 3: Proportion of times an object is predicted in the other language. In parenthesis the number of examples corresponding to the percentage. In bold when $o_{s}^{l_{t}}$ or $o_{t}^{l_{s}}$ are detected more often for each of the experiments. The total number of examples varies, see total numbers in Table 2.
E.1 Different Relation, Different Subject
Figure 20: Probability of the patch answer and the context answer when patching at different layers. Left is XGLM, right mT5.
E.2 Same Relation, Different Subject
Figure 21: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers. Left is XGLM, right mT5.
Figure 22: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers in XGLM, for examples with $\{\neq\mathcal{L},=r,\neq s\}$.
Figure 23: Probability of the $\mathcal{L}_{c}(o_{c})$ when patching at different layers in XGLM, for examples with $\{\neq\mathcal{L},=r,\neq s\}$. Note that the plots do not share the y-axis.
Figure 24: Probability of the $\mathcal{L}_{p}(o_{p})$ when patching at different layers in mT5, for examples with $\{\neq\mathcal{L},=r,\neq s\}$.
Figure 25: Probability of the $\mathcal{L}_{c}(o_{c})$ when patching at different layers in mT5, for examples with $\{\neq\mathcal{L},=r,\neq s\}$. Note that the plots do not share the y-axis.
E.3 Different Relation, Same Subject
Figure 26: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers for examples with $\{\neq\mathcal{L},\neq r,=s\}$. Top is XGLM, bottom mT5.
Figure 27: Probability of the $\mathcal{L}_{p}(o_{p})$ when patching at different layers in XGLM, for examples with $\{\neq\mathcal{L},\neq r,=s\}$.
Figure 28: Probability of the $\mathcal{L}_{c}(o_{c})$ when patching at different layers in XGLM, for examples with $\{\neq\mathcal{L},\neq r,=s\}$. Note that the plots do not share the y-axis.
Figure 29: Probability of the $\mathcal{L}_{p}(o_{p})$ when patching at different layers in mT5, , for examples with $\{\neq\mathcal{L},\neq r,=s\}$.
Figure 30: Probability of the $\mathcal{L}_{c}(o_{c})$ when patching at different layers in mT5, for examples with $\{\neq\mathcal{L},\neq r,=s\}$. Note that the plots do not share the y-axis.
Figure 31: Number of times the patch object is predicted in the context language for the experiment of same relation different subject. Left XGLM, right mT5.
Figure 32: Number of times the object is predicted in the opposite language in mT5 in the different relation same subject experiment.
Figure 33: Number of times the context object is predicted in the patch language in XGLM in the experiment of different relation same subject.

### v2

@@
-How Do Multilingual Models Remember? Investigating Multilingual Factual Recall Mechanisms
+How Do Multilingual Language Models Remember Facts?
 Abstract
 Large Language Models (LLMs) store and retrieve vast amounts of factual knowledge acquired during pre-training.
-Prior research has localized and identified mechanisms behind knowledge recall; however, it has primarily focused on English monolingual models.
-The question of how these processes generalize to other languages and multilingual LLMs remains unexplored.
-In this paper, we address this gap by conducting a comprehensive analysis of two highly multilingual LLMs.
-We assess the extent to which previously identified components and mechanisms of factual recall in English apply to a multilingual context.
-Then, we examine when language plays a role in the recall process, uncovering evidence of language-independent and language-dependent mechanisms.
-How Do Multilingual Models Remember?
-Investigating Multilingual Factual Recall Mechanisms
-Constanza Fierro   Negar Foroutan   Desmond Elliott   Anders Søgaard University of Copenhagen    EPFL
+Prior research has localized and identified mechanisms behind knowledge recall; however, it has only focused on English monolingual models.
+The question of how these mechanisms generalize to non-English languages and multilingual LLMs remains unexplored.
+In this paper, we address this gap by conducting a comprehensive analysis of three multilingual LLMs.
+First, we show that previously identified recall mechanisms in English largely apply to multilingual contexts, with nuances based on language and architecture.
+Next, through patching intermediate representations, we localize the role of language during recall, finding that subject enrichment is language-independent, while object extraction is language-dependent.
+Additionally, we discover that the last token representation acts as a Function Vector (FV), encoding both the language of the query and the content to be extracted from the subject.
+Furthermore, in decoder-only LLMs, FVs compose these two pieces of information in two separate stages.
+These insights reveal unique mechanisms in multilingual LLMs for recalling information, highlighting the need for new methodologies—such as knowledge evaluation, fact editing, and knowledge acquisition—that are specifically tailored for multilingual LLMs.
+How Do Multilingual Language Models Remember Facts?
+Constanza Fierro∗†   Negar Foroutan   Desmond Elliott   Anders Søgaard University of Copenhagen    EPFL
+**footnotetext: Correspondance: Constanza Fierro <c.fierro@di.ku.dk>.
 1 Introduction
-Large Language Models (LLMs) acquire a vast amount of factual knowledge during pre-training, including many propositional facts of the form “The capital of France is __” Petroni et al.
+Large Language Models (LLMs) learn extensive factual knowledge during pre-training, including propositional facts like “The capital of France is __” Petroni et al.
 (2019).
-Such acquisition also occurs in multilingual language models (Kassner et al., 2021; Jiang et al., 2020; Yin et al., 2022), although with considerable variation in performance depending on the input language.
-This raises an important question: Is this variation caused by the knowledge being stored in a language-specific manner, or is it related to the exact phrasing sensitivity?
-(Elazar et al., 2021) If trust turns on knowledge recall (Grasswick, 2010; Hawley, 2012; Nguyen, 2022), such assessments will be integral to establishing whether multilingual models are trustworthy.
-Mechanistic interpretability aims to understand how and why models process information and make predictions at an internal structural level (Elhage et al., 2022; Nanda et al., 2023; Olah et al., 2020).
-Recent works have focused on identifying the specific model components responsible for storing acquired knowledge (Meng et al., 2022; Sharma et al., 2024), and understanding how this knowledge is retrieved internally during inference (Geva et al., 2023; Chughtai et al., 2024).
-However, so far, this research has been exclusively in English and largely limited to autoregressive LMs, except for Sharma et al.
-(2024), who study these mechanisms in Mamba (Gu and Dao, 2023), a state-space LM.
-The acquisition of factual knowledge is usually studied by assessing whether a model predicts the expected fact, given a subject and relation as input, e.g.
-“[Y], which is the capital of France”.
-This setup was originally proposed for masked LMs Petroni et al.
-(2019), where the object to be predicted could be positioned anywhere within the input.
-With the increasing popularity of decoder-only models (Brown et al., 2020), however, the input format has shifted.
-For these models, the placeholder (e.g., “[Y]”) has to be placed at the end of the sentence, so that the model predicts the next token, which is the object (Meng et al., 2022; Yu et al., 2023; Geva et al., 2023; Hase et al., 2023; Wang et al., 2024; Cohen et al., 2024).
-Clearly, this approach presents challenges for languages that do not follow a subject-verb-object (SVO) word order.
-In languages like Turkish, Japanese, Korean, and Persian, which predominantly use a subject-object-verb (SOV) structure, it can be unnatural and difficult to craft inputs that fit this template for studying acquired factual knowledge.
-In this paper, we study the mechanisms of factual recall in multilingual LLMs, focusing on two architectures: the decoder-only model XGLM and the encoder-decoder model mT5; and centering on four key questions: (1) Are the factual recall mechanisms and localization found for English LMs also present in multilingual LLMs?
-(2) Do these mechanisms function the same way for different languages?
-(3) Does the model architecture—encoder-decoder or decoder-only—affect these mechanisms?
-(4) When language is encoded, are there language-dependent mechanisms at play?
-To address (1)–(4), we first conduct a causal tracing analysis (§3) to determine whether the early MLP modules processing the final subject token are as decisive in multilingual models as in English-centric models (Meng et al., 2022).
-In XGLM, this effect is absent, while in mT5, we observe a strong causal impact from both the first and last subject tokens across all layers of the encoder.
-Second, we analyze the mechanisms described in Geva et al.
-(2023), with a focus on when information from the subject and relation flows to the last token, and when the object is extracted from the subject representation (§4).
-We find that information flows similarly in monolingual and multilingual decoder-only models, while for mT5, there are discrepancies.
-In terms of the extraction event, in XGLM, both feed-forward and attention sublayers contribute to it, in contrast to English autoregressive models, where attention modules do most work.
-In mT5, this mechanism is primarily performed by the cross-attention, occurring later in the process and in a more localized manner compared to decoder-only models.
-Finally, we use activation patching (Zhang and Nanda, 2024; Ghandeharioun et al., 2024) and insert the last token representation from an English input into the forward pass of another language (§5); this way, we are able to characterize more fully the multilingual recalling process.
-We find that in XGLM there are two distinct phases.
-First the relation representation propagates to the last token, and then the output language.
-While for mT5, both the relation and language propagate in the later layers, with the relation being encoded slightly earlier.
-Additionally, we show that the subject and relation representations are language-independent, as they can be interchanged and used in different languages; while the extraction event is language-specific.
-The findings of this work contribute to the understanding of crosslingual factual knowledge transfer and consistency, and can further inform knowledge editing methodologies, and trustworthiness assessments in multilingual models.
+While multilingual models also acquire such knowledge, their performance varies significantly across languages (Kassner et al., 2021; Jiang et al., 2020; Yin et al., 2022), raising questions about whether this variation stems from language-specific storage or phrasing sensitivity (Elazar et al., 2021).
+Such assessments will be critical for determining the trustworthiness of multilingual LLMs, if trust relies on knowledge (Grasswick, 2010; Hawley, 2012; Nguyen, 2022).
+Mechanistic interpretability research has begun uncovering how models store and retrieve knowledge internally, with recent studies identifying specific components for knowledge storage (Meng et al., 2022; Sharma et al., 2024) and retrieval mechanisms (Geva et al., 2023; Chughtai et al., 2024).
+However, they have focused exclusively on English LLMs, mainly autoregressive ones.
+[^1] Except Sharma et al.
+(2024), who extend these analyses to Mamba (Gu and Dao, 2023), a state-space language model.
+Encoder-decoder architectures, which could enable better cross-lingual representations (Li et al., 2024), remain underexplored.
+Additionally, recent studies have shown that LLMs share circuits across languages for specific tasks (Ferrando and Costa-jussà, 2024; Zhang et al., 2025), but these are limited to syntactic tasks and do not address how concepts are represented or recalled cross-lingually.
+While Dumas et al.
+(2024) examined how disentangled language is from concepts using a translation task, here we analyze fact recall, which allows us to offer broader insights into how language is encoded.
+In this paper, we study the mechanisms of factual recall in multilingual LLMs, focusing on two architectures: decoder-only (XGLM and EuroLLM) and encoder-decoder (mT5).
+We analyze a simple form of information extraction, where the input contains a subject and a relation, and the model predicts the corresponding object (e.g.
+“Paris” in the earlier example).
+Our analysis centers on three key questions: (1) Does the localization of factual knowledge in English LLMs extend to multilingual LLMs?
+(2) Are the factual recall mechanisms found for English LLMs also present in multilingual LLMs?
+(3) When does language play a role in the recall mechanism?
+To address the first question, we use causal tracing analysis to assess if early MLP modules processing the final subject token are as decisive in multilingual models as in English-centric models (Meng et al., 2022).
+Our results (§3) show that EuroLLM and mT5 exhibit strong causal effects for all subject tokens, with EuroLLM exhibiting this in earlier layers and mT5 across all encoder layers.
+In XGLM, the last subject token has a stronger effect, but with two early sites, unlike English GPT’s single site.
+Additionally, in all models, MLPs in later layers are decisive in recovering factual information, a contrast to Meng et al.
+(2022) findings.
+Next, we investigate the second question by analyzing the mechanisms described in Geva et al.
+(2023); namely, the three-step process, where the relation information flows to the last token, then the subject flows, and finally the attention layers extract the object (§4).
+We find that in multilingual decoder-only models the information flows similarly to monolingual models, but differently for the encoder-decoder model mT5.
+In terms of the extraction event, in decoder-only LLMs, both feed-forward and attention sublayers contribute to it, unlike English autoregressive models, where attention modules dominate.
+In mT5, this mechanism is primarily performed by the cross-attention.
+Overall, our findings indicate that some of the localization (§3) and mechanisms (§4) of fact retrieval in English LLMs generalize to multilingual LLMs, but with key variations.
+Finally, to address the third question, we investigate where language plays a role within the three-step process described by Geva et al.
+(2023), to characterize how and where factual knowledge cross-lingual transfer may occur.
+Using activation patching (Zhang and Nanda, 2024; Ghandeharioun et al., 2024) we insert the intermediate representation of the last token from an English forward pass into the forward pass of another language (§5).
+Our results reveal that the last token acts as a Function Vector (FV) Todd et al.
+(2024), encoding both the relation and the output language, which is then applied to the subject in the context.
+Crucially, the fact that the FV formed in one language can be used with an input in another language demonstrates that the subject and relation representations are largely language-independent, while the extraction event is language-specific.
+Furthermore, in decoder-only LLMs (XGLM, EuroLLM), the FV is constructed in two distinct phases: first, it encodes only the relation, and later, the language is incorporated.
+These findings advance our understanding of factual recall, with implications for cross-lingual knowledge transfer, knowledge editing, and trustworthiness in multilingual LLMs.
+Our results on language flow open new avenues for studying whether models ‘think’ in English (Wendler et al., 2024), by examining how the FV is altered across languages.
+Additionally, the late-site causal effect of MLPs and their dominance in the extraction phase suggest that fact editing techniques and knowledge evaluations must extend beyond early MLPs and attention layers (Tamayo et al., 2024).
 2 Experimental Setup
-We focus on the basic form of factual knowledge recall, where LMs are tasked with predicting the correct object $o$ for a given subject $s$ and a relation $r$.
-These $(s,r,o)$ triplets are obtained from WikiData, and natural language templates are used to describe the relation, with placeholders for the subject and object.
-For the relation born-in, a template could, e.g., be “[X] was born in [Y]”, where [X] represents the placeholder for the subject and [Y] is the object to be predicted by the model.
-We use the mParaRel dataset (Fierro and Søgaard, 2022), with triplets and templates for 45 languages.
-[^1] We augment mParaRel by fetching WikiData aliases of the objects to be able to identify more cases of correct object prediction.
-We filter out trivial examples where the object is directly mentioned in the query, e.g., South Asia belongs to the continent of.
-For our analysis, we select 10 typologically diverse languages representing a variety of scripts, language families, and word orders, including SVO, SOV, and VSO structures: English (en), Spanish (es), Vietnamese (vi), Turkish (tr), Russian (ru), Ukrainian (uk), Japanese (ja), Korean (ko), Hebrew (he), Persian (fa), and Arabic (ar).
-We analyze two different architectures of multilingual language models, XGLM (Lin et al., 2021),
-Language XGLM mT5
-en 1812 1543
-es 1380 1192
-vi 1646 993
-tr 418 1058
-ru 830 683
-uk 213 456
-ko 116 630
-ja 6 358
-he 13 565
-fa 7 406
-ar 811 488
-Table 1: Number of memorized knowledge triplets $(s,r,o)$.
-a decoder-only architecture with 7.5B parameters and 32 layers, and mT5-xl (Xue et al., 2021), an encoder-decoder architecture of 3.7B parameters and 24/24 encoder/decoder layers.
-For XGLM, we only use the templates in mParaRel that have the object placeholder at the end of the sentence, while for mT5, we use all the templates available.
-XGLM pre-training data includes 30 languages but it does not include uk, he, and fa; while mT5 includes 101 languages and covers all the ones studied here.
-As we aim to investigate the process of knowledge recall, and similarly to previous work (Meng et al., 2022; Geva et al., 2023), we only consider examples where the model predicts the correct object completion.
-Since mParaRel provides multiple templates (paraphrases) for each relation, we greedy-decode for every available template corresponding to a given triplet and check for an exact match.
-[^2] We perform greedy decoding of a maximum of 50 tokens and check for exact match of any of the object aliases.
-In practice, the 50 tokens generation is an overestimate, as on average, around 0.5 tokens are generated prior to the target object (for some templates an article or other filling tokens are generated before predicting the object).
-If multiple templates trigger the correct object, we randomly select one for the analyses.
-Once an exact match is found, any extra tokens generated before the object are added to the example input text.
-This ensures that when we feed again the example to the model, the next token predicted is the first token of the object.
-[^3] For mT5, we feed the mParaRel input into the encoder with a sentinel token in the object placeholder.
-In the decoder, we provide the beginning-of-sequence token followed by the sentinel token.
-We only check the tokens generated next to the first sentinel token, as the pre-training task of the model is to generate the text for each sentinel in the input, the decoder usually continues to generate answers for other sentinel tokens.
-Any tokens generated before the target object, are added to the decoder tokens since adding these to the encoder does not ensure that the next token predicted will be the object.
+We focus on a simple form of factual knowledge recall, where LMs are tasked with predicting the correct object $o$ for a given subject $s$ and a relation $r$.
+These $(s,r,o)$ triplets are obtained from WikiData, and natural language templates are used to describe the relation, with placeholders for the subject and object.222 For example, the relation born-in could use the template “[X] was born in [Y]”, where [X] is the subject and [Y] is the object to be predicted.
+For our analysis, we select 10 typologically diverse languages representing various scripts, families, and word orders: English (en), Spanish (es), Vietnamese (vi), Turkish (tr), Russian (ru), Ukrainian (uk), Japanese (ja), Korean (ko), Hebrew (he), Persian (fa), and Arabic (ar).
+See Table 2 for language characteristics.
+Models
+We analyze decoder-only and encoder-decoder architectures.
+[^3] For decoder-only models, we only use the templates in mParaRel that have the object placeholder at the end of the sentence, while for encoder-decoder, we use all the templates.
+The decoder-only LLMs are XGLM (Lin et al., 2021),
+XGLM EuroLLM mT5
+en 1812 2332 1543
+es 1380 1913 1192
+vi 1646 779 993
+tr 418 799 1058
+ru 830 1680 683
+uk 213 1244 456
+ko 116 308 630
+ja 6 42 358
+he 13 107 565
+fa 7 31 406
+ar 811 1790 488
+Table 1: Number of facts $(s,r,o)$ correctly predicted.
+with 7.5B parameters and 32 layers, and EuroLLM (Martins et al., 2024) with 9B parameters and 42 layers; the encoder-decoder mT5-xl (Xue et al., 2021) has with 3.7B parameters and 24 encoder-decoder layers.
+The pre-training data varies across models: mT5 is pretrained on 101 languages, covering all the languages in our study; XGLM covers 30 languages, excluding uk, he, and fa; while EuroLLM covers 35 languages, excluding vi, he, and fa.
+Data
+We use the mParaRel dataset (Fierro and Søgaard, 2022), which includes triplets and templates for 45 languages.
+These templates are machine translations of the ParaRel English templates (Elazar et al., 2021).
+[^4] We augment the objects in mParaRel, filter out trivial examples, and when there are enough examples, we use a crosslingual subset (see Appendix B).
+To investigate the process of knowledge recall we only consider examples where the model predicts the correct object completion (Meng et al., 2022; Geva et al., 2023).
+Since mParaRel provides multiple paraphrased templates for each relation, we greedy-decode for every available template corresponding to a given triplet and check for an exact match.
+If multiple templates yield a match, we randomly select one for the analysis.
+In cases where an article or other filler tokens precede the object, we include these tokens in the input text to ensure that when the example is fed into the model for our analysis, the next predicted token is the first token of the object (Implementation details in Appendix B.1).
 Table 1 presents the number of examples for which the correct object is predicted.
-Due to the limited number of examples memorized by XGLM in ja, he, and fa, we exclude these languages from our main analysis of XGLM.
+We exclude languages with too few examples from our analysis, namely ko, ja, he, and fa for XGLM, and ja, he, and fa for EuroLLM.
+Notation
+Given a transformer model with $L$ layers, let $h_{t}^{l}$ be the representation of the token $t$ at layer $l$.
+When the model is an encoder-decoder, let $e_{i}^{l}$ be the representation of the encoder layer $l$ for the $i$-th token in the encoder input.
+Then, the encoder layer computes $h_{t}^{l+1}=h_{t}^{l}+s^{l}+f^{l}$ and the decoder $h_{t}^{l+1}=h_{t}^{l}+s^{l}+c^{l}+f^{l}$, where $s^{l}=\text{Self Attn.}(h_{0}^{l}\ldots h_{t}^{l})$, $c^{l}=\text{Cross Attn.}(h_{t}^{l},e_{0}^{l}\ldots e_{n}^{l})$ and $f^{l}=\text{MLP}(h_{t}^{l}+s^{l}+c^{l})$.
+If decoder-only, then $c^{l}$ does not apply.
 3 Causal Tracing
@@
 (2022), we trace the causal effects of hidden states using causal mediation analysis (Pearl, 2022).
-To do this, we feed the model a corrupted input and selectively restore a particular component in the computational graph, such as the attention output of layer $l$.
-We then measure how much of the original probability (from the uncorrupted input) is recovered, thereby assessing the indirect influence of the restored component on the model’s prediction.
-Specifically, we corrupt the input by adding Gaussian noise to the subject tokens, [^4] We follow Meng et al.
+Let $\mathbb{P}(o)$ be the probability of the predicted object token, and $LS(o)$ its logit score.
+We corrupt the input by adding Gaussian noise to the subject tokens, [^5] We follow Meng et al.
 (2022) and add $\epsilon\sim\mathcal{N}(0,(3\sigma)^{2})$, with $\sigma$ being the standard deviation of the subjects tokens embeddings from the data used.
-and observe the probability $\hat{\mathbb{P}}$ of the originally predicted token.
-Then, we run inference again on the corrupted input, but this time, we restore a specific hidden state in the model and track the probability ^$\hat{\mathbb{P}}_{\text{restored}}$ of the originally predicted token.
-We study the indirect effect $\text{IE}=\hat{\mathbb{P}}_{\text{restored}}-\hat{\mathbb{P}}$.
-Consider a transformer model with $L$ layers; let $h_{t}^{l}$ be the representation of the token $t$ at layer $l$.
-When the model is an encoder-decoder, let $e_{i}^{l}$ be the representation of the encoder layer $l$ for the $i$-th token in the input.
-Then, the encoder layer computes $h_{t}^{l+1}=h_{t}^{l}+s^{l}+f^{l}$ and the decoder $h_{t}^{l+1}=h_{t}^{l}+s^{l}+c^{l}+f^{l}$, where $s^{l}=\text{Self Attn.}(h_{0}^{l}\ldots h_{t}^{l})$, $c^{l}=\text{Cross Attn.}(h_{t}^{l},e_{0}^{l}\ldots e_{n}^{l})$ and $f^{l}=\text{MLP}(h_{t}^{l}+s^{l}+c^{l})$.
-If decoder-only, then $c^{l}$ does not apply.
-Let $h_{t}^{l}$ be the hidden state from the clean run and $\overline{h}_{t}^{l}$ that of the corrupted run.
-We restore a state by setting $\overline{h}_{t}^{l}\leftarrow h_{t}^{l}$, and similarly, we restore the self-attention layers contribution by setting $\overline{s}^{l}\leftarrow s^{l}$ for all the attention modules in a window of size $w$ (analogous for $c$ and $f$).
+and observe the corrupted probability ~$\tilde{\mathbb{P}}(o)$ of the originally predicted token.
+Then, we run inference again on the corrupted input, but this time, we restore a specific hidden state in the model and track the probability ~$\tilde{\mathbb{P}}_{\text{restored}}(o)$.
+We study the indirect effect of such component as $\text{IE}_{\mathbb{P}}=\tilde{\mathbb{P}}_{\text{restored}}(o)-\tilde{\mathbb{P}}(o)$, or if using logits $\text{IE}_{LS}=\tilde{LS}(o)-\tilde{LS}_{\text{restored}}(o)$.
+Specifically, we restore a state by setting $\tilde{h}_{t}^{l}\leftarrow h_{t}^{l}$, where $h_{t}^{l}$ is the hidden state from the clean run and $\tilde{h}_{t}^{l}$ that of the corrupted run; and similarly, we restore the self-attention layers contribution by setting $\tilde{s}^{l}\leftarrow s^{l}$ for all the attention modules in a window of size $w$ (analogous for $c$ and $f$).
 [^6] We use windows because, generally, the contributions of the sub-layers are gradual (Geva et al., 2021).
 In other words, multiple layers contribute the same behavior to the residual, and their sum produces the observed effect.
-Following Meng et al.
-(2022), we set $w=10$ for XGLM, and $w=6$ for mT5 (as it has 24 layers).
-We repeat the corrupted run with restoration ten times, each with different noise samples, and report the average.
+We set $w$=$4$ for XGLM, $w$=$5$ for EuroLLM, and $w$=$3$ for mT5, to restore $\sim 12\%$ layers.
+We repeat the corrupted run with restoration ten times with different noise samples, and report the average.
 We compare our results in multilingual LLMs to those of Meng et al.
-(2022), who analyzed the factual recall of GPT-2 XL in English, [^6] They used the CounterFact dataset, which is based on the ParaRel dataset.
-Thus, the English data is the same as the one we use here.
-and reached two key conclusions.
-First, they identified an “early site”, where the MLPs processing the last subject token in the early and middle layers play a crucial role in recovering from input corruption.
-Second, they found a “late site”, where the attention modules processing the last token in the later layers also significantly contribute to prediction recovery.
+(2022), who analyzed the factual recall of GPT-2 XL in English [^7] They used the CounterFact dataset, which is based on ParaRel.
+Thus, their English data is the same as our data.
+and reached two key conclusions: (1) they identified an “early site”, where the MLPs processing the last subject token in the early and middle layers play a crucial role in recovering from input corruption; and (2), they found a “late site”, where the attention modules processing the last token in the later layers also significantly contribute to prediction recovery.
 [^8] While they consider the late site unsurprising, as it directly precedes the final prediction, this observation primarily applies to hidden state restoration, not necessarily attention restoration.
@@
 (2023) referred to as the extraction event (§4).
-We present the causal analysis plots for each model and language in Appendix B and present our main observations here.
-We do not find for XGLM an early site of MLP modules nor for English or any of the other languages.
-Thus, we conclude that this feature is specific to English-centric models, as Sharma et al.
-(2024) also observed it for another English-centric architecture (Mamba).
-We do observe a causal effect of the last subject token when restoring the state in the early layers.
-This is somewhat expected as it reverts the corruption performed in the input, but it is important to note until which layer it is possible to recover from this corruption only by restoring the last subject token.
-It also highlights the relevance of the last subject token in decoder-only models for all languages.
-Figure 1: Attention knockout between the last token and a given set of tokens. Each layer represents the effect of the knockout on a window of $w$ layers. Top XGLM ($w=6$), bottom mT5 ($w=4$).
-For mT5, we do observe a causal effect from MLP modules of the subject tokens, but it is not an early site as in GPT.
-Rather it is a causal effect along all the layers of the encoder, in both the first and last token of the subject, with some languages having a stronger effect in the last subject token.
-In terms of late site, we find it in the self attention for XGLM, and in the cross attention layers in mT5.
-For both mT5 and XGLM, we also observe a late site of the MLP at the last token, though this effect is smaller than that of the attention or the subject.
-We conclude that the multilingual models behave in a similar way for the different languages, finding that the causal effects of both MLP and attention modules are positioned the same, and that for decoder-models the last subject token (at the early layers) is critical for the recalling, while for encoder-decoder models the MLPs processing both the first and last subject token encode the necessary information (at any encoder layer).
+Figure 1: Attention knockout between the last token and a given set of tokens.
+Each layer represents the effect of the knockout on a window of $w$ layers.
+Top EuroLLM ($w=7$), bottom mT5 ($w=4$).
+XGLM in Figure 21.
+We present the causal analysis plots for each model and language in Appendix C and present our main observations here.
+[^9] In line with Zhang and Nanda (2024) we find that analyzing the $\text{IE}_{\mathbb{P}}$ overestimates the causal effect of some tokens over others.
+For example, in XGLM by looking at the $\text{IE}_{\mathbb{P}}$ we would not find an early site, which is present when studying the $\text{IE}_{LS}$ (see Figure 11 vs Figure 9), or in EuroLLM the last subject token would seem more relevant than the other subject tokens (see Figure 15 vs Figure 13).
+So we base our main observations using $\text{IE}_{LS}$.
+Our results indicate that the last subject token does not consistently have a stronger effect on information recovery in early MLPs.
+In EuroLLM and mT5, all subject tokens exhibit similar effects, while in XGLM, the last subject token has a stronger influence.
+In EuroLLM, the early site spans half of the initial layers, whereas in mT5, it covers all encoder layers.
+For XGLM, the early site occurs twice: once in the initial layers and again in the middle layers.
+Regarding the late site, we observe a strong causal effect in MLP layers before the very last layers in all models.
+The layers where these sites occur are consistent across languages, with only a slight variation in the late site’s endpoint for EuroLLM.
+These findings suggest that some, but not all, conclusions from English LLMs generalize to multilingual LLMs.
+The MLP early site extends across models, and the MLP late site, which was not observed by Meng et al.
+(2022), was present in all models studied.
+This has implications for knowledge localization and fact editing methodologies, which should account for both early and late MLPs in multilingual contexts.
 4 Factual Recall Components
@@
 Information Flow
-Like Geva et al.
-(2023), we use attention knockout to visualize the propagation of information to the last token.
-Based on insights from the causal analysis, we intervene in the last token self-attention in XGLM and cross-attention for mT5.
-Specifically, we set the attention score to be zero between the last token and another set of tokens $\{t\}$.
-Given that the contributions of attention blocks are incremental, we perform this intervention within a window of size $w$ around layer $l$.
-The set of tokens $\{t\}$ consists of (1) subject tokens, (2) non-subject tokens (relation tokens), and (3) the last token (in mT5, the last token refers to the last token in the decoder input).
-We use a window of size 6 for XGLM and of size 4 for mT5, following Geva et al.
-(2023) to knock out around 18% of layers.
-We then measure the change in probability of the originally predicted first token (without knockout), aggregating the results across examples by calculating the relative difference (i.e., the change in probability proportional to the original probability).
-Thus, a big drop in relative probability indicates a critical flow of information at that layer, from the selected tokens to the last token.
-We present the results in Figure 1 (plots per language in Appendix C).
-For XGLM model, we find broadly similar dynamics to those of GPT (Geva et al., 2023) with a critical flow from the subject at later layers around 25, and negligible flow from the last token to itself.
-However, there is a strong propagation of information from the non-subject tokens throughout the early and middle layers, shown by the drop of -80% in probability, which reaches -100% for some languages in layers 5-10.888GPT2-xl attention to non-subject tokens is more pronounced in the early layers (-40%), decreasing from there.
-GPT-J attends to non-subject tokens through early and mid layers but with a weaker drop (-20%) (Geva et al., 2023).
-Most languages follow a similar pattern, but ru benefits from the knockout to the subject token in early layers; for ar and ko, XGLM attends strongly to non-subject tokens until the end.
-The mT5 results show a different behavior.
-First, the last token does attend to itself, a pattern that is more pronounced in the early and later layers, but less so around the middle layers.
-[^9] We hypothesize that this is probably due to the last token encoding information about which sentinel token is generated at this step (indicating where in the input it should fill), since we see an almost identical curve for all the languages when the last token cannot attend to itself or when it cannot attend to the sentinel token in the decoder (Figure 14).
-Second, the last token attends to non-subject tokens across all layers, with a weaker drop at the early layers compared to XGLM.
-Finally, similar to XGLM, there is critical information flowing from the subject token in the later layers.
-However, this drop is larger and occurs even later in the computation compared to XGLM.
-Most languages behave similarly, with the exception of ja and ko, exhibiting less flow from last and non-subject tokens in the early layers.
-Figure 2: Extraction rates. Left XGLM, right mT5.
+[…]
+Figure 1 presents the results with plots per language in Appendix D.
+Firstly, the results show that in each model the information flows fairly similarly for all the languages, and similar patterns to those found in English GPT are present (Geva et al., 2023).
+On the one hand, the subject information flows to the last token most critically at the later layers in all models.
+On the other hand, the relation flows throughout all the layers, with EuroLLM and mT5 having more similar curves.
+Nevertheless, we note two differences from the conclusions reached with English GPT (Geva et al., 2023): (1) the propagation of non-subject tokens to the last token does not strictly precedes the subject propagation, in XGLM the drop is the highest in the earlier layers but it continues to be important until the end, and in mT5 and EuroLLM this information has two peaks, in the early and later layers; and (2) the last token of mT5 encodes critical information that flows from one layer to the next through the cross-attention (as opposed to the negligible flow of the last token in the decoder-only models).
+[^10] We hypothesize this may occur because the last token encodes which sentinel token is being generated, indicating where to fill in the input.
+Since we see an almost identical curve when the last token cannot attend to the sentinel token in the decoder (Figure 24).
+Figure 2: Extraction rates. Left XGLM, right EuroLLM, bottom mT5.
 Prediction Extraction
-Like Geva et al.
-(2023), we measure extraction events at each layer of the model.
-Using the notation from §3, let $h^{l}$ be the representation of the last token at layer $l$, and let $E$ be the embedding matrix, then the predicted token is $o^{*}=\arg\,\max(Eh^{L})$.
-We say there is an extraction event at the self-attention at layer $l$ if $\arg\,\max(Es^{l})=o^{*}$ (similarly for $c^{l}$ and $f^{l}$).
+Following Geva et al.
+(2023), we measure extraction events at each layer.
+Let $h^{l}$ be the representation of the last token at layer $l$, and let $E$ be the embedding matrix; the predicted token is $o=\arg\,\max(Eh^{L})$.
+An extraction event occurs at layer $l$ if $\arg\,\max(Es^{l})=o$ (similarly for $c^{l}$ and $f^{l}$).
 The extraction rate is the proportion of examples for which an extraction event occurs at a given layer.
-Figure 2 shows the extraction rates for both models, with a detailed language breakdown in Appendix D.
-A key difference between XGLM and GPT-2 (Geva et al., 2023) is the prominence of MLP modules in the extraction of the object in XGLM.
-We note that this is not simply because the MLP is forwarding the already extracted object from the preceding attention layer.
-We plot in Figure 17 how often an MLP has an extraction event while the preceding attention has not, and we find a similar plot to the one in Figure 2, where the MLP has a higher peak in layer 25 than the attention.
-Additionally, we observe that for English (Figure 15), in line with the results for GPT-2, the attentions have higher extraction rate than the MLPs.
-For mT5, the cross-attention layers are the main mechanism for extraction, although the MLP layers also contribute to a lesser degree, resembling GPT-2’s behavior.
-However, for certain languages (he, fa, ko, and ru), the MLPs show comparable or even higher extraction rates (Figure 16).
-Finally, we observe that in mT5, extraction events occur later in the computation, with peaks at layers 19 and 22.
-5 Information Flow Through Patching
-Finally, we study the role of language in the recall prediction of transformer models.
-We employ activation patching (Wang et al., 2022) by intervening in the model’s computation to identify critical steps.
-Through these analyses, we contribute three key elements to the understanding of the recall process in transformer models: (1) we localize exactly at which layer the relation information flows to the last token; (2) we identify the extraction mechanism in multilingual LLMs; and (3) we explore how language plays a role in recall prediction.
-Figure 3: Patching strategy.
-The patch example (left) and the context example (right).
-The patch of this transformer is performed at layer 0 and it affects the computation of layer 1, thus the output is perturbed.
-Let $h^{l}$ be the representation at layer $l$ of the last token in the input.
-For a given input $p$, we take $h^{l}_{p}$ and patch it into the forward pass of another input $c$ at the same layer; that is, we set $h^{l}_{t,c}\leftarrow h^{l}_{t,p}$, and then continue the forward pass.
+Figure 2 shows the extraction rates for each model, with a detailed language breakdown in Appendix E.
+Our results demonstrate that extraction events can be detected in multilingual LLMs, though rates vary across languages.
+A key finding is the prominence of MLP modules in object extraction for multilingual decoder-only models (XGLM and EuroLLM).
+To rule out the possibility that MLPs simply forward extracted objects from preceding attention layers, we measure how often MLPs perform an extraction without prior attention extraction (Figure 28).
+We find that MLPs indeed perform the extraction for most languages, though in English, attention modules dominate, aligning with GPT-2 results (Geva et al., 2023).
+In mT5, cross-attention layers drive the extraction, with MLPs playing a secondary role, resembling GPT-2’s behavior.
+Additionally, in mT5 the extraction events occur in later layers, with peaks in 19 and 22.
+The results show that in multilingual LLMs, information flow from non-subject tokens to the final token is less localized compared to English LLMs, possibly because the model processes relational and language information at different stages.
+Additionally, the extraction mechanism in multilingual LLMs is more complex, involving both attention and MLP modules.
+This implies that editing techniques should focus not only on early MLPs enhancing subject representations but also on the later MLPs that extract the object.
+5 Language Information Flow
+Previous analyses do not provide insights into how the language is included and used during recall.
+In this section, we use activation patching, similarly to Dumas et al.
+(2024), to: (1) disentangle when the relation information flows to the last token and when the language information flows, (2) identify which tokens contribute the language information, and (3) interpret the last token representation as a Function Vector Todd et al.
+(2024) [^11] Function Vectors (FV) were originally defined as the mean vector of outputs from specific attention heads.
+Here, we interpret the last token representation as an FV because it triggers a specific execution for different contexts..
+Figure 3: Patching strategy. The patch example’s last token is inserted in the context example’s forward pass, which perturbs the output of subsequent layers.
+(a) Counts of examples where the patch produces the object prediction.
+(b) Histogram of patches that cause the context answer to be predicted in the patch language, $\mathcal{L}_{p}(o_{c})$.
+Figure 5: Prediction tokens from (a) $\{=\mathcal{L},\neq r,\neq s\}$ and (b) $\{\neq\mathcal{L},=r,\neq s\}$ for XGLM, EuroLLM, mT5.
+Let $h^{l}_{t}$ be the representation at layer $l$ of the last token in the input.
+For a given input $p$, we take $h^{l}_{t,p}$ and patch it into the forward pass of another input $c$ at the same layer; that is, we set $h^{l}_{t,c}\leftarrow h^{l}_{t,p}$, and then continue the forward pass.
 We refer to the example $p$ as the patch example, and $c$ as the context example (see Figure 3).
 By patching the last token representation in each layer and measuring the model’s predicted output, we can study when the representation contains the information about relation, language, and predicted object, and how cross-lingual these representations are.
-An input example expresses a relation $r$ and a subject $s$ in language $\mathcal{L}$.
+An input example expresses a relation $r$ and a subject $s$ in language $\mathcal{L}$, which we note $(\mathcal{L}(r),\mathcal{L}(s))$.
 We conduct three experiments where the patch and context examples share certain input characteristics.
-In particular, we test three different conditions: (1) $\{=\mathcal{L},\neq r,\neq s\}$, where the language is the same but the relation and subject differ; (2) $\{\neq\mathcal{L},=r,\neq s\}$, where the relation remains the same but the language and subject differ; and (3) $\{\neq\mathcal{L},\neq r,=s\}$, where the subject is the same but the language and relation differ.
+In particular, we test: (1) $\{=\mathcal{L},\neq r,\neq s\}$, where the language is the same but the relation and subject differ; (2) $\{\neq\mathcal{L},=r,\neq s\}$, where the relation remains the same but the language and subject differ; and (3) $\{\neq\mathcal{L},\neq r,=s\}$, where the subject is the same but the language and relation differ.
 (1) allows us to study how the relation is encoded while controlling for the language, whereas (2)-(3) help us explore how the interaction between relation and language affects the extraction of the object from the subject.
 In all experiments, English is used as the patch language, $\mathcal{L}_{p}=\text{en}$.
-Let the input of the patch be $(\mathcal{L}_{p}(r_{p}),\mathcal{L}_{p}(s_{p}))$ and the input of the context be $(\mathcal{L}_{c}(r_{c}),\mathcal{L}_{c}(s_{c}))$.
-When there is no intervention, the model correctly predicts $\mathcal{L}_{p}(o_{r_{p},s_{p}})$ and $\mathcal{L}_{c}(o_{r_{c},s_{c}})$ respectively.
+Let the patch input be $(\mathcal{L}_{p}(r_{p}),\mathcal{L}_{p}(s_{p}))$ and the context input be $(\mathcal{L}_{c}(r_{c}),\mathcal{L}_{c}(s_{c}))$.
+Without intervention, the model correctly predicts $\mathcal{L}_{p}(o_{r_{p},s_{p}})$ and $\mathcal{L}_{c}(o_{r_{c},s_{c}})$ respectively.
 [^12] For simplicity, we may refer to the object as $\mathcal{L}_{p}(o_{p})$ or $\mathcal{L}_{c}(o_{c})$ when $r$ and $s$ are from the same input.
-To analyze the effect of patching, we study both the output probabilities and the final predictions.
-First, we evaluate the probability of the originally predicted token, $\mathcal{L}_{p}(o_{p})$ and $\mathcal{L}_{c}(o_{c})$, and similar to §4 we aggregate across examples by calculating the relative difference (the change in probability proportional to the original (unpatched) probability).
-Second, we examine the predicted token, and we check whether it corresponds to $\mathcal{L}_{p}(o_{c})$, the context answer but in the patch language, or to $\mathcal{L}_{c}(o_{p})$, the patch answer but in the context language.
-Note that we consider only pairs of patch-context examples for which we have the corresponding objects, and which do not share the same spelling.
-[^11] We limit ourselves to pairs where: (1) mParaRel has the expected object in the other language (e.g.
-that we have the aliases of the context object in the patch language $\mathcal{L}_{p}(o_{c})$); and (2) the prediction can only match one of the object aliases sets, e.g.
-to say that the predicted token $t=\mathcal{L}_{p}(o_{c})$ then $t\neq\mathcal{L}_{c}(o_{c})$, which may happen if the languages share the spelling of the object $o_{c}$.
-For example, “Asia” is spelled the same in English and Spanish.
-As a result, the number of examples varies when we study output probabilities, predictions of $\mathcal{L}_{c}(o_{p})$, or predictions of $\mathcal{L}_{p}(o_{c})$.
-See Table 2 for the number of examples for each model and experiment.
+To analyze patching effects, we examine both output probabilities and predicted tokens.
+For probabilities, we aggregate across examples by calculating the change relative to the original (unpatched) probability, as in §4.
+For predicted tokens, we check if they match the patch or context object, or a variant with swapped language or relation (e.g., if $\mathcal{L}_{p}(o_{c})$ is predicted in the setups (2)-(3)).
+We only consider valid patch-context pairs where $o_{r_{p},s_{c}}$ and $o_{r_{c},s_{p}}$ exist, and for predicted token analysis, we enforce distinct spellings.
+[^13] E.g.
+to claim the predicted token is $t=\mathcal{L}_{p}(o_{c})$ we require $t\neq\mathcal{L}_{c}(o_{c})$, as languages may share spellings (e.g., “Asia” in English and Spanish).
+Consequently, the number of examples varies across analyses of output probabilities and token predictions.
+See Table 4 and 6 for the number of examples for each model and experiment.
+Figure 6: Relative probability change of $\mathcal{L}_{c}(o_{c})$ when patching $\{\neq\mathcal{L},=r,\neq s\}$, for XGLM, EuroLLM, mT5.
+Figure 7: Relative probability change of $\mathcal{L}_{c}(o_{c})$ when patching $\{\neq\mathcal{L},\neq r,=s\}$, for XGLM, EuroLLM, mT5.
 Different Relation, Different Subject
 First, we analyze the pairs of examples with $\{=\mathcal{L},\neq r,\neq s\}$.
-[^12] For example, $r_{p},s_{p}=$“The capital of France is” and $r_{c},s_{c}=$“The language spoken in Germany is”.
+For example, $r_{p},s_{p}=$“The capital of France is” and $r_{c},s_{c}=$“The language spoken in Germany is”.
 Then, $o_{r_{p},s_{c}}=$ Berlin and $o_{r_{c},s_{p}}=$ French.
-We study $\sim$1000 examples for which mParaRel has the objects $o_{r_{c},s_{p}}$ and $o_{r_{p},s_{c}}$.
-[^13] The total number of examples that have different spellings when checking for $o_{r_{p},s_{c}}$ is 984 (994) in XGLM (mT5), and for $o_{r_{c},s_{p}}$ is 974 (990).
-We present the prediction results in Figure 4, with corresponding probability plots in Figure 20.
-We observe that the relation information is propagated to the last token representation by layer 13 in XGLM and layer 16 in mT5, since the patching at that point starts producing the object corresponding to the relation in the patch $o_{r_{p},s_{c}}$.
-By layer 22 in XGLM and layer 18 in mT5, we observe that the object has been fully extracted and is encoded within the last token representation, as predictions from these layers correspond to $o_{r_{p},s_{p}}$.
-Layer 22 in XGLM aligns with the first attention peak in the extraction rate of English (Figure 15), and layer 18 in mT5 matches the first cross-attention peak (Figure 16).
-This shows that the extraction measured by the vocabulary projection indeed represents the point at which the object is extracted and encoded in the last token representation.
-If the object was encoded earlier and only decoded with the extraction event, we would see $o_{r_{p},s_{p}}$ being predicted from patches at earlier layers.
-Figure 4: Counts of examples where the patch produces the object prediction. Top is XGLM, bottom mT5.
+We sample 1000 examples for which mParaRel has the objects $o_{r_{c},s_{p}}$ and $o_{r_{p},s_{c}}$.
+We present the prediction results in Figure 5(a), and probability plots in Figure 32.
+The relation information is contained in the last token representation when the green curve increases, since patching at that point starts producing the object corresponding to the relation in the patch $o_{r_{p},s_{c}}$.
+When the green curve decreases and the blue curve increases, the object has been fully extracted and is encoded in the last token vector, as predictions from these layers correspond to $o_{r_{p},s_{p}}$ and no information is taken from the context $c$.
+The layers where this happens align with the peaks of the extraction rate in English (Figure 25-27).
+This shows that the extraction measured by the vocabulary projection represents the point at which the object is extracted and encoded in the last token representation.
+If the object was encoded earlier and only decoded with the extraction event,$o_{r_{p},s_{p}}$ would be predicted from patches at earlier layers.
 Same Relation, Different Subject
@@
 Is the language information entangled to the subject or the relation representation?
-Or is the language encoded separately from these two?
-Here, we study pairs of examples with $\{\neq\mathcal{L},=r,\neq s\}$.
-[^14] For example: “France’s capital is” and “La capital de Alemania es” (Gloss: “The capital of Germany is”).
-If we obtain $\mathcal{L}_{p}(o_{c})$, this suggests that the language is encoded in the last token representation before the extraction happens.
-On the other hand, if we obtain $\mathcal{L}_{c}(o_{p})$, we could infer that the language is encoded after the extraction.151515 After filtering examples with similar object spellings, the total number of pairs we have per language is around 150-800.
-We present the probability of the context answer token $\mathcal{L}_{c}(o_{c})$ in Figure 5 (the probability of $\mathcal{L}_{p}(o_{p})$ and per language plots are provided in Appendix E.2).
-First, we observe that the probability of $\mathcal{L}_{p}(o_{p})$ increases sharply around layers 21-25 for XGLM and around 18-21 for mT5.
-For XGLM, regarding the probability of $\mathcal{L}_{c}(o_{c})$, we observe that patching in the early layers generally hurts the model’s performance for most languages.
-However, patching in the middle layers (10-20) either increases the probability or results in only a minimal decrease.
-Given that the relation is the same in both examples, this suggests that by these middle layers, $r$ is encoded in $h^{l}_{p}$, and in the subsequent layers, the subject and language of the context are integrated to yield the final prediction, $\mathcal{L}_{p}(r)+\mathcal{L}_{c}+s_{c}=\mathcal{L}_{c}(o_{c})$.
+Here, we study pairs of patch-context with $\{\neq\mathcal{L},=r,\neq s\}$, e.g., “France’s capital is” and “La capital de Alemania es” (Gloss: “The capital of Germany is”).
+If we obtain $\mathcal{L}_{p}(o_{c})$ (“Berlin”), it would suggest that the language is encoded in the last token representation before the extraction happens.
+On the other hand, if we obtain $\mathcal{L}_{c}(o_{p})$ (“París”), we could infer that the language is encoded after the extraction.
+We present the probability of the context answer token $\mathcal{L}_{c}(o_{c})$ in Figure 7.141414The probability of $\mathcal{L}_{p}(o_{p})$ and per language plots are provided in Appendix F.2).
+We observe that for decoder-only models, patching in the early layers generally hurts the model’s performance for most languages, while, patching in the middle layers either increases the probability or results in only a minimal decrease.
+Given that the relation is the same in both examples, this suggests that by these middle layers, $r$ is encoded in $h^{l}_{p}$, and in the subsequent layers the subject and language of the context are integrated to yield the final prediction, $\mathcal{L}_{p}(r)+\mathcal{L}_{c}+s_{c}=\mathcal{L}_{c}(o_{c})$.
 Moreover, for some languages the relation representation from the patch is better than the one constructed using the context input, as the relative probability is positive.
-Figure 5: Relative probability change of $\mathcal{L}_{c}(o_{c})$ when patching the last token representation between pairs of examples with the same relation and a different subject (Top XGLM, bottom mT5).
-In the case of mT5, however, the probability of $\mathcal{L}_{c}(o_{c})$ consistently decreases, plateauing in the middle layers before dropping to zero by layer 22.
+In the case of mT5, however, the probability of $\mathcal{L}_{c}(o_{c})$ consistently decreases, plateauing in the middle layers before dropping to zero.
 From the attention knockout analysis, we recall that in mT5, the subject is integrated into the last token representation only after layer 15, while the relation is consistently represented throughout the middle layers.
-Therefore, these patching results imply that the relation encoded in the last token from the patch input does not help in retrieving the correct context object in the target language in mT5, whereas it proves useful in XGLM.
-Figure 6: Histogram of patches that cause the context answer to be predicted in the patch language, $\mathcal{L}_{p}(o_{c})$ (Top XGLM, bottom mT5).
-In terms of predicted tokens, we find that across-the-board $\mathcal{L}_{c}(o_{p})$ is rarely predicted while $\mathcal{L}_{p}(o_{c})$ appears frequently (Table 3).
-This suggests that the last token representation encodes the patch language $\mathcal{L}_{p}$ but not yet the object, as the object is derived from the subject context.
-Figure 6 shows the layers where $\mathcal{L}_{p}(o_{c})$ is predicted (plot for $\mathcal{L}_{p}(o_{c})$ in Figure 31).
-As discussed in §4, the extraction event typically happens around layers 18-28 in XGLM and 16-22 in mT5.
-Therefore, we observe that the language is attended to at the beginning of the extraction phase, because if we patch earlier, the output is in the context language.
-Then, the language is not attended anymore but rather the language information that already flowed to the last token is the one used for the rest of the extraction phase.
-We can interpret the last token representation as containing the “task” (Todd et al., 2024), the relation that needs to be extracted and in which language, which is used in the extraction event.
-This task vector can be transferred to contexts in another language in both models, as the task vector is constructed from the patch input and used in the context subject representation to predict $\mathcal{L}_{p}(o_{c})$.
+Therefore, these patching results imply that the relation encoded in the last token from the patch input does not help in retrieving the correct context object in the context language in mT5, whereas it proves useful in XGLM and EuroLLM.
+In terms of predicted tokens, we find across-the-board that $\mathcal{L}_{p}(o_{c})$ is frequently predicted while $\mathcal{L}_{c}(o_{p})$ is not (Table 5).
+This suggests that the last token representation encodes the patch language $\mathcal{L}_{p}$ but not yet the object, as the object is derived from the context subject.
+We plot the layers where $\mathcal{L}_{p}(o_{c})$ is predicted in Figure 5(b) ($\mathcal{L}_{c}(o_{p})$ in Figure 47).
+We can conclude that the language information flows to the last token right before the peak of $\mathcal{L}_{p}(o_{c})$ predictions, because if we patch earlier, the output is in the context language (see the probabilities of $\mathcal{L}_{c}(o_{c})$ in Figure 7).
+Moreover, these peaks match the beginning of their corresponding extraction phases, thus the language information flows right before the extraction phase.
+As a result, we interpret the last token representation as containing a Function Vector (FV), the relation that needs to be extracted and in which language, which is used in the extraction event.
+The FV can be transferred to contexts in another language, as the FV is constructed from the patch input and is used in the context subject representation to predict $\mathcal{L}_{p}(o_{c})$.
 Different Relation, Same Subject
-We just saw that the representation from the patch will encode at some point the output language but not yet the object.
-In the previous experiment both examples had the same relation, so it could be that we observe the prediction $\mathcal{L}_{p}(o_{r,s_{c}})$ because the relation is language specific and encodes the output language.
-To analyze if this is the case, we now apply patching on examples with different languages and relations but the same subject $\{\neq\mathcal{L},\neq r,=s\}$.
-[^16] For example: “[Y]’s capital is” and “El idioma oficial de [Y] es” (Gloss: “The official language of [Y] is”).
-The number of examples is around 150-300 per language, while for XGLM in uk and tr is around 30-70 and only 6 in ko (Table 2).
-When analyzing the predictions of $\mathcal{L}_{c}(o_{p})$ and $\mathcal{L}_{p}(o_{c})$, as we restrict the examples to have different spellings, the total number of pairs is around 75-250 depending on the language.
-Figure 7: Relative probability change of $\mathcal{L}_{c}(o_{c})$ when patching the last token representation between pairs of examples with the different relation and the same subject (Top XGLM, bottom mT5).
-We present the probability of the original token $\mathcal{L}_{c}(o_{c})$ and $\mathcal{L}_{p}(o_{p})$ in Figures  7 and 26, respectively.
-We observe that the probability of $\mathcal{L}_{p}(o_{p})$ starts increasing in layers 10-15 for es, vi, tr, and ko while the increase starts from layers 15-20 for ru, uk, and ar.
-In mT5, this increase is more uniform across languages and occurs slightly later.
-As for $\mathcal{L}_{c}(o_{c})$, unlike the previous experiment, we observe that in XGLM, the probability decreases early for all languages (except tr), plateaus around the middle layers, and then drops to zero by the mid-layer range.
-In mT5, similar to the previous experiment, the probability of $\mathcal{L}_{c}(o_{c})$ drops at the beginning but, instead of plateauing, it continues to decline until it reaches zero around layer 17.
-In terms of predictions, we find that $\mathcal{L}_{c}(o_{p})$ is frequently predicted for XGLM (Figure 8), while $\mathcal{L}_{p}(o_{c})$ appears but less often (see Figure 31 and Table 3).
-This suggests that in layers 15-21, the patched last token representation primarily captures the relation $r_{p}$, without yet encoding the output language or subject information.
-Therefore, in XGLM, the relation and language representations are disentangled, as the relation flows to the last token before the output language does.
+We just saw that for all models the representation from the patch will encode at some point the output language but not yet the object.
+It could be that we observed the prediction $\mathcal{L}_{p}(o_{c})$ because the relation is language specific and encodes the output language.
+To analyze if this is the case, we now apply patching on examples with different languages and relations but the same subject $\{\neq\mathcal{L},\neq r,=s\}$, e.g., “France’s capital is” and “El idioma oficial de Francia es” (Gloss: “The official language of France is”).
+We observe that the probability of $\mathcal{L}_{c}(o_{c})$ (Figures  7) in decoder-only models, unlike the previous experiment, decreases early for all languages (except tr and ko), plateaus around the middle layers, and then drops to zero by the mid-layer range.
+For mT5, the probability drops at the beginning but, instead of plateauing as before, it continues to decline until it reaches zero.
+When compared to the former experiment (Figure 7), this suggests that the relation information is encoded in the middle layers, as $\mathcal{L}_{c}(o_{c})$ decreases earlier when the patch and context have different relations.
+In terms of predictions, we find that $\mathcal{L}_{c}(o_{p})$ is frequently predicted for XGLM and EuroLLM (Figure 8), while $\mathcal{L}_{p}(o_{c})$ appears but less often (Table 7).
+In line with the previous observation, the plot shows that the last token representation in the middle layers where $\mathcal{L}_{c}(o_{p})$ is predicted, primarily captures the relation from the patch $r_{p}$, without yet encoding the output language or subject information (as these are taken from the context).
+Therefore, in decoder-only models, the relation and language representations are disentangled, as the relation flows to the last token before the output language does.
 Allowing the relation to be combined with different languages.
-As for mT5, we observe very few examples where $\mathcal{L}_{c}(o_{p})$ or $\mathcal{L}_{p}(o_{c})$ are predicted, which aligns with the findings of the two former experiments, where we see that the relation is encoded in the last token in layers 15-21 (Figure 4), and the language flows to the last token around layer 15-19 (Figure 6).
+As for mT5, we observe very few examples where $\mathcal{L}_{c}(o_{p})$ or $\mathcal{L}_{p}(o_{c})$ are predicted, which aligns with the findings of the two former experiments, where we see that the relation is encoded in the last token in layers 15-21 (Figure 5(a)), and the language flows to the last token around layer 15-19 (Figure 5(b)).
 We conclude that both the relation and language flow to the last token around the same time, and thus, in this experiment, we cannot see a disentangled behavior.
-This presents an interesting contrast with XGLM.
+This presents an interesting contrast with decoder-only models.
 The decoder in mT5 has access to the same encoder representations throughout all its layers, so it does not need (and thus does not learn) to attend to these earlier or in different stages.
-By contrast, in a decoder-only model like XGLM, the last token has access to representations that evolve across layers, so it learns to attend to relevant information when it becomes most salient.
+By contrast, in a decoder-only model, the last token has access to representations that evolve across layers, so it learns to attend to relevant information when it becomes most salient.
 Nonetheless, we cannot reach a definite conclusion on whether the language representation in mT5 is entangled or not to the relation representation.
 More detailed analysis of what is being attended when the relation flows and when the language flows should be performed in future work.
-Figure 8: Histogram of patches that cause the prediction to be $\mathcal{L}_{c}(o_{p})$.
+Figure 8: Patches that cause the prediction to be $\mathcal{L}_{c}(o_{p})$. Top XGLM, bottom EuroLLM.
 6 Conclusion
-In this paper, we analyzed the mechanisms involved in the recall of factual knowledge in 10 languages using multilingual transformer-based LMs.
-We compare these findings to previous research on English recall in English autoregressive LLMs.
-We discovered that some mechanisms, such as the flow of subject representations in the later layers and the extraction phase, are shared between multilingual and monolingual models.
-However, we also identified notable differences, including the absence of knowledge localization in the early MLPs and the joint role of MLPs and attention modules during the extraction phase in multilingual models.
-Finally, we also studied when language encoding occurs during the recall process by patching representations.
-In XGLM, we observed two distinct phases: initially, the relation flows to the last token, followed by the language.
-In contrast, in mT5, both the relation and language flow to the last token at similar layers.
-This suggests that while the relation and subject representations are multilingual and can be used to extract objects across languages, the extraction phase itself appears to be language-specific, as the language is encoded prior to extraction.
-These findings provide new evidence to understand the factual knowledge recall in transformer LMs, contributing with analyzes in different architectures and languages, and suggesting new directions for the study of cross-lingual transfer.
+In this paper, we analyzed factual knowledge recall mechanisms in 10 languages using multilingual transformer-based LMs, comparing them to prior research on English recall in monolingual autoregressive LLMs.
+We discovered that some mechanisms, such as the flow of subject representations in the later layers and the extraction phase, are present in multilingual and monolingual models.
+However, we also identified notable differences, including the localization of knowledge in not only one but two MLP sites; and the joint role of MLPs and attention modules during the extraction phase in multilingual models.
+A key contribution of our work is the first-ever investigation of language encoding during recall, achieved through patching representations.
+In decoder-only models, the relation flows to the last token first, followed by the language.
+In contrast, in mT5, both relation and language flow to the last token at similar layers.
+This suggests that while relation and subject representations are multilingual and enable cross-lingual object extraction, the extraction phase itself is language-specific, as language encoding precedes extraction.
+These findings provide new evidence to understand the factual knowledge recall in transformer LMs, and to how decoder-only LMs resolve tasks in stages.
+Contributing with new directions for the study of cross-lingual transfer and knowledge localization.
 7 Limitations
-In this paper, we examined two model architectures, leaving out the effects of model sizes, instruction fine-tuning, or models like Llama that can behave multilingually but have less coverage and less multilingual pre-training data.
+In this paper, we examined three model architectures, leaving out the effects of model sizes, instruction fine-tuning, or models like Llama that can behave multilingually but have less coverage and less multilingual pre-training data.
 Additionally, our analyses were conducted on 500-1000 examples per language, which we believe provides a sufficient sample size for generalization; however, the results are inherently limited by the relations present in the mParaRel dataset, which may not capture all factual nuances.
 Additionally, although we analyzed 10 diverse languages, many more languages exist, and further research is needed to confirm the generalizability of our findings across a broader linguistic spectrum.
-Lastly, we described the main mechanisms found in XGLM and mT5, however other weaker mechanisms could be at play, which could describe, for example, the low extraction rates found for some languages (Figure 16) or the few examples where the object seems to be encoded from early layers before the extraction takes place (Figure 31).
+Lastly, we described the main mechanisms found in XGLM, EuroLLM and mT5, however other weaker mechanisms could be at play, which could describe, for example, the low extraction rates found for some languages (Figure 27) or the few examples where the object seems to be encoded from early layers before the extraction takes place (Figure 47).
 Appendix A Related Work
-Factual Knowledge Recall.
+Factual Knowledge Recall
 Petroni et al.
-(2019) analyzed the factual knowledge encoded in pre-trained language models, leading to further multilingual research.
+(2019) analyzed factual knowledge in pre-trained language models, by creating the LAMA dataset, which contains natural language templates and WikiData triplets.
+Subsequently, Elazar et al.
+(2021) developed the ParaRel dataset, which contains manually paraphrases of the LAMA templates, to measure the consistency of factual knowledge.
 [^15] One may object to calling the ability to solve cloze tests as knowledge.
@@
 (2024) in thinking knowledge may nevertheless be the appropriate concept, since LLMs are mostly consistent and can sometimes provide justification, e.g., in virtue of world models or training data attribution.
-Kassner et al.
+This led to further research in multilingual settings, Kassner et al.
 (2021); Jiang et al.
-(2020) extended this work by translating the LAMA dataset (Petroni et al., 2019) and evaluating the factual knowledge acquired by multilingual pre-trained models, while Wang et al.
-(2024) evaluated it from the perspective of knowledge edits (Zhu et al., 2021).
+(2020) translated LAMA, creating mLAMA and X-FACTR, and evaluated the factual knowledge acquired by multilingual pre-trained models, while Wang et al.
+(2024a) evaluated knowledge from the perspective of knowledge edits and their impact between Chinese and English (Zhu et al., 2021).
 Building on this, Yin et al.
 (2022) curated a dataset that focuses on the cultural aspect of factual knowledge.
-While Fierro and Søgaard (2022) presented the mParaRel dataset181818mParaRel is a machine-translated version of the English templates from ParaRel (Elazar et al., 2021), where subjects and objects are translated using WikiData identifiers.
-to examine the consistency of factual predictions in multilingual pre-trained language models.
+While Fierro and Søgaard (2022) studied multilingual consistency of knowledge using mParaRel, a machine-translated version of the English templates in ParaRel, where subjects and objects are translated using WikiData identifiers.
 In a related effort, Qi et al.
 (2023) measured the cross-lingual consistency of multilingual LLMs by evaluating the consistency across languages, rather than just across paraphrases in the same language.
-Interpretability on Factual Knowledge Recall.
+Qi et al.
+(2023) used the mLAMA dataset.
+In this paper, we used mParaRel since it is an extended version of mLAMA, as it is a translation of the paraphrases in mParaRel.
+Interpretability on Factual Knowledge Recall
 As discussed earlier, Meng et al.
@@
 (2024) examined whether a linear transformation of the subject is sufficient to represent various relations, analyzing GPT and Llama models with English data.
-Activation Patching.
+Activation Patching
 Ghandeharioun et al.
@@
 In their experiments, they focused on “decoding” the contents of a latent representation by patching it into the forward pass of another input.
-Similarly, Zhang and Nanda (2024) explored the differences and implementation nuances of patching representations versus adding Gaussian noise.
-More relevant to the type of patching we conducted in this work, Dumas et al.
-(2024) investigated a word translation task to assess whether the model encodes concepts that are disentangled from the language during translation.
-This builds on earlier findings by Wendler et al.
-(2024), who observed that models like Llama-2 (Touvron et al., 2023) seem to use English as an internal pivot language.
-Consistent with our results, Dumas et al.
-(2024); Foroutan et al.
-(2022) provided evidence for the existence of language-agnostic representations.
+More closely related to our work, Dumas et al.
+(2024) patched the last token representation to investigate if models encode language-agnostic concepts in a word translation task.
+Their findings, consistent with ours and those of Foroutan et al.
+(2022), support the existence of language-agnostic representations.
+Similarly, Wang et al.
+(2024b) patched the last token representation to study how relations are encoded in factual English queries, finding that the last token encodes the query’s relation at a specific computational stage—aligning with our results.
+In this paper, we are able to draw more general conclusions given that we study a cross-lingual factual recall setup.
+While Dumas et al.
+(2024) concluded that LLMs first resolve the output language, we find that the model first resolves the query’s relation and then the language.
+Thus, we interpret the last token representation as a Function Vector (FV) (Todd et al., 2024), encoding the function to solve the given task: for Dumas et al.
+(2024) the task was translation so the function encodes the language to which translate.
+Finally, we are also able to show that for decoder-only models the FV composes the language onto the relation FV, which extends the findings from Wang et al.
+(2024b).
 Multilingual Language Models
-In this work, we focused on generative models and selected mT5 (Xue et al., 2021) and XGLM Lin et al.
-(2021) due to their broad language coverage, highly multilingual training data, and pre-training with up-sampling of low-resource languages to ensure more balanced language representation.
+In this work, we focused on generative models and selected mT5 (Xue et al., 2021), XGLM Lin et al.
+(2021), and EuroLLM (Martins et al., 2024) due to their broad language coverage, highly multilingual training data, and pre-training with up-sampling of low-resource languages to ensure more balanced language representation.
 However, many other models are available.
@@
 Llama-2 (Touvron et al., 2023) has been used in some multilingual analyses (Wendler et al., 2024; Dumas et al., 2024), but it is primarily an English-centric model given that 90% of its pre-training data is in English.
-More recent models that could have been considered include mGPT (Shliazhko et al., 2022) and EuroLLM (Martins et al., 2024), both decoder-only models trained on 61 and 34 languages, respectively.
+More recent models that could have been considered include mGPT (Shliazhko et al., 2022), a decoder-only model trained on 61.
 On the other hand, the Aya-101 model (Üstün et al., 2024) is an instruction fine-tuned version of mT5; however, we focus on pre-trained multilingual LMs to compare our analysis to the English monolingual pre-trained versions previously studied.
-Appendix B Causal Tracing
+Language Script Family SOV/SVO
+English Latin Germanic SVO
+Spanish Latin Romance SVO
+Vietnamese Latin Austroasiatic SVO
+Turkish Latin Turkic SOV
+Russian Cyrillic Slavic SVO*
+Ukrainian Cyrillic Slavic SVO*
+Japanese Kanji Proto-Japonic SOV
+Korean Korean Koreanic SOV
+Hebrew Hebrew Arabic VSO
+Farsi (Persian) Perso-Arabic Indo-Iranian SOV
+Arabic Arabic Arabic VSO
+Table 2: Languages, their scripts, families, and sentence structures (SVO: subject-verb-object, SOV: subject-object-verb, VSO: verb-subject-object, SVO*: SVO dominant but SOV is also possible).
+Appendix B Experimental Setup
+We use the mParaRel triplets and templates to query the factual knowledge of the language models.
+As mentioned earlier, we perform 3 modifications to the dataset.
+First, we fetch WikiData for aliases of the target object to be able to match different possible surface forms.
+Second, we filter out examples where the target object is contained in the query, e.g.
+Microsoft Outlook is developed by.
+Finally, to better compare across languages we control the variety of the subject-object pairs, by only using a crosslingual version of mParaRel.
+Specifically, for each relation we filter out triplets that are not present in all the languages (in mParaRel a subject and object may have not been translated if they were not found in WikiData).
+Thus, in the crosslingual mParaRel version, each subject-object pair in a relation is present in each of the languages.
+For XGLM we additionally restrict the templates to have the object at the end of the sentence, therefore for some languages (tr, ko, ja, and fa) both the autorregressive condition on the templates and the crosslingual restriction leads to too few total examples (< 1000), so for these we do not impose the crosslingual restriction.
+For all the other languages, and for all the languages in mT5 there are enough examples so we restrict them to be crosslingual.
+In Table 3 we present the total number of examples we consider per language and model, and the correctly predicted fraction.
+[^16] In decoder-only models, the total number of examples is higher in some languages due to crosslingual filtering, which is applied per relation.
+If a language has no examples for a given relation, it doesn’t restrict examples in other languages.
+Since decoder-only models use only autoregressive templates, some lower-resource languages may have zero examples for certain relations.
+However, when using all templates (as in mT5), these lower-resource languages can restrict the triplets available for other languages within the same relation.
+XGLM and EuroLLM mT5
+Language Correct Total Percentage Correct Total Percentage
+en 1812 4147 43.7% 1543 3853 40.0%
+es 1380 4167 33.1% 1192 3926 30.4%
+vi 1646 4068 40.5% 993 3748 26.5%
+tr 418 2278 18.3% 1058 4033 26.2%
+ru 830 4133 20.1% 683 3826 17.9%
+uk 213 4144 5.1% 456 3830 11.9%
+ko 116 1444 8.0% 630 3783 16.7%
+ja 6 1790 0.3% 358 4124 8.7%
+he 13 4403 0.3% 565 4064 13.9%
+fa 7 4388 0.2% 406 3814 10.6%
+ar 811 4482 18.1% 488 4110 11.9%
+Table 3: Total number of examples in mParaRel.
+B.1 Prompt Details
+For mT5, we feed the mParaRel input into the encoder with a sentinel token in the object placeholder.
+In the decoder, we provide the beginning-of-sequence token followed by the sentinel token.
+We only check the tokens generated next to the first sentinel token, as the pre-training task of the model is to generate the text for each sentinel in the input, the decoder usually continues to generate answers for other sentinel tokens.
+Any tokens generated preceding the object, are added to the decoder input since adding these to the encoder does not ensure that the next token predicted will be the object.
+B.2 Computational Resources
+The experiments were run in A100 GPUs. The causal analysis took from 30 minutes to 12 hours depending on the language. The rest of the experiments took 1-2 hours per language.
+Appendix C Causal Tracing
+C.1 XGLM
+Figure 9: XGLM causal analysis for each language (continues in Figure 10).
+Average logit score (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 4 layers, or the Self Attn.
+in a window of 4 layers.
 Figure 10: XGLM causal analysis for each language (Rest of the languages in Figure 9).
-Average probability (of the originally predicted token) recovered when corrupting the subject in the input and restoring the hidden representation at a given layer (State), or restoring the MLP / Self Attn.
-in a window of 10 layers.
+Average logit score (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 4 layers, or the Self Attn.
+in a window of 4 layers.
+Figure 11: XGLM causal analysis for each language (continues in Figure 12).
+Average probability (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 4 layers, or the Self Attn.
+in a window of 4 layers.
 Figure 12: XGLM causal analysis for each language (Rest of the languages in Figure 11).
-Average probability (of the originally predicted token) recovered when corrupting the subject in the input and restoring the hidden representation at a given layer (State), or restoring the MLP / Self Attn.
-in a window of 10 layers.
-Figure 11: Causal analysis of mT5 for each language (Rest of the languages in Figure 12).
-Average probability (of the originally predicted token) recovered when corrupting the subject in the input and restoring the hidden representation at a given layer (State), or restoring the MLP / Self Attn.
-/ Cross Attn.
-in a window of 6 layers.
-Figure 12: Causal analysis of mT5 for each language (Rest of the languages in Figure 11).
-Average probability (of the originally predicted token) recovered when corrupting the subject in the input and restoring the hidden representation at a given layer (State), or restoring the MLP / Self Attn.
-/ Cross Attn.
-in a window of 6 layers.
-Appendix C Attention Knockout
+Average probability (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 4 layers, or the Self Attn.
+in a window of 4 layers.
+C.2 EuroLLM
+Figure 13: EuroLLM causal analysis for each language (continues in Figure 14).
+Average logit score (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 5 layers, or the Self Attn.
+in a window of 5 layers.
+Figure 14: EuroLLM causal analysis for each language (Rest of the languages in Figure 13).
+Average logit score (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 5 layers, or the Self Attn.
+in a window of 5 layers.
+Figure 15: EuroLLM causal analysis for each language (continues in Figure 16).
+Average probability (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 5 layers, or the Self Attn.
+in a window of 5 layers.
+Figure 16: mT5 causal analysis for each language (Rest of the languages in Figure 15).
+Average probability (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 5 layers, or the Self Attn.
+in a window of 5 layers.
+C.3 mT5
+Figure 17: mT5 causal analysis for each language (continues in Figure 18).
+Average logit score (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 3 layers, or the Self Attn.
+in a window of 3 layers.
+Figure 18: mT5 causal analysis for each language (Rest of the languages in Figure 17).
+Average logit score (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 3 layers, or the Self Attn.
+in a window of 3 layers.
+Figure 19: mT5 causal analysis for each language (continues in Figure 20).
+Average probability (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 3 layers, or the Self Attn.
+in a window of 3 layers.
+Figure 20: mT5 causal analysis for each language (Rest of the languages in Figure 19).
+Average probability (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 3 layers, or the Self Attn.
+in a window of 3 layers.
+Appendix D Attention Knockout
+[Uncaptioned image][Uncaptioned image][Uncaptioned image]
+Figure 21: Attention knockout between the last token and a given set of tokens.
+Each layer represents the effect of the knockout on a window of $w$ layers.
+Models from top to bottom: XGLM ($w=6$), EuroLLM ($w=7$), mT5 ($w=4$).
 Figure 22: XGLM attention knockout for each language.
+Figure 23: EuroLLM attention knockout for each language.
 Figure 24: Attention knockout for each language in mT5.
-We present the results from knocking out the information from the last token in the decoder (last) and the masked object in the encoder (enc_sentinel).
-However, from the encoder sentinel there is no much flow of information so these are not plot in the main body of the paper.
-Appendix D Extraction Event
+The knockout is performed between the last token in the decoder (“last”) to a given set of tokens $\{t\}$, and between the masked token in the encoder (“enc_sentinel”) and $\{t\}$.
+From the encoder sentinel there is no much flow of information so these were not included in the main body.
+Appendix E Extraction Event
 Figure 25: XGLM extraction rates for each language.
+Figure 26: EuroLLM extraction rates for each language.
 Figure 27: mT5 extraction rates for each language.
-Figure 17: Number of extraction events split by precedence (or not) of an extraction event in the self-attn (cross-attn). Left XGLM, right mT5.
+Figure 28: Number of extraction events split by precedence (or not) of an extraction event in the self-attn or cross-attn. Models from top to bottom: XGLM, EuroLLM, mT5.
 Figure 29: Number of extraction events split by precedence (or not) of an extraction event in the self-attn in XGLM for each language.
+Figure 30: Number of extraction events split by precedence (or not) of an extraction event in the self-attn in EuroLLM for each language.
 Figure 31: Number of extraction events split by precedence (or not) of an extraction event in the cross-attn in mT5 for each language.
-Appendix E Patching
-XGLM mT5
-$=$ Relation, $\neq$ Subject $\neq$ Relation, = Subject $=$ Relation, $\neq$ Subject $\neq$ Relation, = Subject
-Source - Target All $o_{s}^{l_{s}}\neq o_{s}^{l_{t}}$ $o_{t}^{l_{s}}\neq o_{t}^{l_{t}}$ All $o_{s}^{l_{s}}\neq o_{s}^{l_{t}}$ $o_{t}^{l_{s}}\neq o_{t}^{l_{t}}$ All $o_{s}^{l_{s}}\neq o_{s}^{l_{t}}$ $o_{t}^{l_{s}}\neq o_{t}^{l_{t}}$ All $o_{s}^{l_{s}}\neq o_{s}^{l_{t}}$ $o_{t}^{l_{s}}\neq o_{t}^{l_{t}}$
-en-es 849 155 140 325 75 82 842 830 779 263 250 249
-en-vi 862 369 377 337 177 180 854 844 823 209 207 204
-en-tr 974 260 253 68 36 15 815 802 794 209 206 205
-en-ru 823 565 588 218 187 177 796 795 787 191 191 190
-en-uk 914 734 715 35 32 32 853 850 852 136 136 133
-en-ko 995 993 991 6 6 6 744 744 743 165 165 165
-en-he 962 903 458 0 0 0 778 778 778 166 166 166
-en-fa 537 528 513 0 0 0 714 714 712 114 114 114
-en-ar 829 818 816 212 209 211 775 775 774 159 159 159
-Table 2: Total number of pairs of examples considered in the patching experiments.
-XGLM mT5
-$=$ Relation, $\neq$ Subject $\neq$ Relation, $=$ Subject $=$ Relation, $\neq$ Subject $\neq$ Relation, $=$ Subject
-Patch - Context $\mathcal{L}_{c}(o_{r,s_{p}})$ $\mathcal{L}_{p}(o_{r,s_{c}})$ $\mathcal{L}_{c}(o_{r_{p},s})$ $\mathcal{L}_{p}(o_{r_{c},s})$ $\mathcal{L}_{c}(o_{r,s_{p}})$ $\mathcal{L}_{p}(o_{r,s_{c}})$ $\mathcal{L}_{c}(o_{r_{p},s})$ $\mathcal{L}_{p}(o_{r_{c},s})$
-en-es 1.3% (2) 38.6% (54) 49.3% (37) 7.3% (6) 0.7% (6) 21.6% (168) 0.8% (2) 10.4% (26)
-en-vi 0.5% (2) 67.1% (253) 11.3% (20) 13.3% (24) 0.2% (2) 20.4% (168) 1.0% (2) 13.7% (28)
-en-tr 1.2% (3) 52.6% (133) 47.2% (17) 33.3% (5) 0.2% (2) 26.3% (209) 2.4% (5) 11.2% (23)
-en-ru 2.3% (13) 53.1% (312) 63.6% (119) 6.2% (11) 1.4% (11) 36.1% (284) 7.9% (15) 11.6% (22)
-en-uk 1.1% (8) 26.2% (187) 15.6% (5) 18.8% (6) 0.4% (3) 31.8% (271) 4.4% (6) 9.0% (12)
-en-ko 0.4% (4) 45.0% (446) 83.3% (5) 33.3% (2) 0.5% (4) 32.3% (240) 4.2% (7) 13.3% (22)
-en-he 0.0% (0) 57.9% (265) - - 0.3% (2) 38.4% (299) 5.4% (9) 16.3% (27)
-en-fa 0.0% (0) 42.3% (217) - - 0.8% (6) 25.3% (180) 3.5% (4) 16.7% (19)
-en-ar 8.8% (72) 39.7% (324) 40.7% (85) 4.7% (10) 0.9% (7) 32.0% (248) 2.5% (4) 9.4% (15)
-Table 3: Proportion of times an object is predicted in the other language.
+Appendix F Patching
+XGLM EuroLLM mT5
+Patch - Context All $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ All $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ All $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$
+en-es 849 834 759 880 862 755 842 831 780
+en-vi 862 857 836 930 910 843 854 844 825
+en-tr 974 961 961 966 945 914 815 803 794
+en-ru 823 821 817 846 846 827 796 795 787
+en-uk 914 914 911 915 915 915 853 850 852
+en-ko 995 995 991 991 991 991 744 744 743
+en-he 962 962 961 810 810 775 778 778 778
+en-fa 537 537 532 771 771 734 714 714 712
+en-ar 829 829 828 862 859 858 775 775 774
+en-ja 134 134 134 895 895 872 774 774 773
+Table 4: Total number of patch-context examples considered in the patching experiments with $\{\neq\mathcal{L},=r,\neq s\}$.
+The $\mathcal{L}_{c}(o_{p})$ column is the total number of examples where the detection of $\mathcal{L}_{c}(o_{p})$ would be unambiguous, that is, $\mathcal{L}_{c}(o_{p})\neq\mathcal{L}_{p}(o_{p})$, conversely for the $\mathcal{L}_{p}(o_{c})$ column.
+XGLM EuroLLM mT5
+Patch - Context $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$
+en-es 0.7% (6) 26.0% (197) 5.0% (43) 34.2% (258) 0.5% (4) 21.5% (168)
+en-vi 0.5% (4) 34.4% (288) 0.5% (5) 16.1% (136) 0.1% (1) 20.2% (167)
+en-tr 0.2% (2) 16.4% (158) 0.2% (2) 25.4% (232) 0.4% (3) 26.2% (208)
+en-ru 2.2% (18) 51.4% (420) 4.6% (39) 63.5% (525) 1.1% (9) 36.1% (284)
+en-uk 0.5% (5) 21.3% (194) 6.9% (63) 62.5% (572) 0.2% (2) 31.8% (271)
+en-ko 0.3% (3) 47.1% (467) 0.1% (1) 75.7% (750) 0.3% (2) 32.3% (240)
+en-he 0.0% (0) 27.3% (262) 0.0% (0) 65.9% (511) 0.4% (3) 38.4% (299)
+en-fa 0.0% (0) 41.7% (222) 0.3% (2) 64.0% (470) 1.0% (7) 25.1% (179)
+en-ar 8.7% (72) 39.5% (327) 4.5% (39) 65.7% (564) 0.6% (5) 31.9% (247)
+en-ja 0.0% (0) 3.0% (4) 1.9% (17) 11.5% (100) 0.0% (0) 26.0% (201)
+Table 5: Proportion of times an object is predicted in the other language in the patching experiments with $\{\neq\mathcal{L},=r,\neq s\}$.
 In parenthesis the number of examples corresponding to the percentage.
-In bold when $o_{s}^{l_{t}}$ or $o_{t}^{l_{s}}$ are detected more often for each of the experiments.
+In bold when $\mathcal{L}_{c}(o_{p})$ or $\mathcal{L}_{p}(o_{c})$ are detected more often for each of the experiments.
 The total number of examples varies, see total numbers in Table 4.
-E.1 Different Relation, Different Subject
-Figure 20: Probability of the patch answer and the context answer when patching at different layers. Left is XGLM, right mT5.
-E.2 Same Relation, Different Subject
-Figure 21: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers. Left is XGLM, right mT5.
+XGLM EuroLLM mT5
+Patch - Context All $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ All $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ All $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$
+en-es 324 284 298 413 360 391 263 253 249
+en-vi 336 329 333 105 102 100 209 207 204
+en-tr 88 77 84 111 102 110 209 207 204
+en-ru 218 216 218 420 415 419 191 191 190
+en-uk 35 35 35 351 347 350 136 136 133
+en-ko 46 46 46 98 98 98 165 165 165
+en-ar 212 212 212 391 390 391 159 159 159
+en-he - - - 19 19 19 166 166 166
+en-ja - - - 1 1 1 85 85 85
+en-fa - - - - - - 114 114 114
+Table 6: Total number of patch-context examples considered in the patching experiments with $\{\neq\mathcal{L},\neq r,=s\}$.
+The $\mathcal{L}_{c}(o_{p})$ column is the total number of examples where the detection of $\mathcal{L}_{c}(o_{p})$ would be unambiguous, that is, $\mathcal{L}_{c}(o_{p})\neq\mathcal{L}_{p}(o_{p})$, conversely for the $\mathcal{L}_{p}(o_{c})$ column.
+XGLM EuroLLM mT5
+Patch - Context $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$
+en-es 18.0% (51) 3.7% (11) 41.7% (150) 2.8% (11) 0.8% (2) 7.6% (19)
+en-vi 9.1% (30) 6.6% (22) 3.9% (4) 8.0% (8) 1.0% (2) 12.3% (25)
+en-tr 36.4% (28) 6.0% (5) 56.9% (58) 4.5% (5) 2.9% (6) 9.3% (19)
+en-ru 64.8% (140) 4.1% (9) 61.7% (256) 9.3% (39) 6.3% (12) 8.9% (17)
+en-uk 17.1% (6) 17.1% (6) 66.0% (229) 2.3% (8) 3.7% (5) 9.0% (12)
+en-ko 41.3% (19) 21.7% (10) 71.4% (70) 5.1% (5) 2.4% (4) 12.1% (20)
+en-ar 40.1% (85) 1.9% (4) 52.8% (206) 0.8% (3) 1.9% (3) 7.5% (12)
+en-he - - 0.0% (0) 0.0% (0) 3.6% (6) 15.7% (26)
+en-ja - - 0.0% (0) 0.0% (0) 0.0% (0) 25.9% (22)
+en-fa - - - - 2.6% (3) 15.8% (18)
+Table 7: Proportion of times an object is predicted in the other language in the patching experiments with $\{\neq\mathcal{L},\neq r,=s\}$.
+In parenthesis the number of examples corresponding to the percentage.
+In bold when $\mathcal{L}_{c}(o_{p})$ or $\mathcal{L}_{p}(o_{c})$ are detected more often for each of the experiments.
+The total number of examples varies, see total numbers in Table 6.
+F.1 Different Relation, Different Subject
+Figure 32: Probability of the patch answer and the context answer when patching at different layers. Models from top to bottom: XGLM, EuroLLM, mT5.
+F.2 Same Relation, Different Subject
+Figure 33: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers. Models from top to bottom: XGLM, EuroLLM, mT5.
 Figure 34: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers in XGLM, for examples with $\{\neq\mathcal{L},=r,\neq s\}$.
 Figure 35: Probability of the $\mathcal{L}_{c}(o_{c})$ when patching at different layers in XGLM, for examples with $\{\neq\mathcal{L},=r,\neq s\}$. Note that the plots do not share the y-axis.
+Figure 36: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers in EuroLLM, for examples with $\{\neq\mathcal{L},=r,\neq s\}$.
+Figure 37: Probability of the $\mathcal{L}_{c}(o_{c})$ when patching at different layers in EuroLLM, for examples with $\{\neq\mathcal{L},=r,\neq s\}$. Note that the plots do not share the y-axis.
 Figure 38: Probability of the $\mathcal{L}_{p}(o_{p})$ when patching at different layers in mT5, for examples with $\{\neq\mathcal{L},=r,\neq s\}$.
 Figure 39: Probability of the $\mathcal{L}_{c}(o_{c})$ when patching at different layers in mT5, for examples with $\{\neq\mathcal{L},=r,\neq s\}$. Note that the plots do not share the y-axis.
-E.3 Different Relation, Same Subject
-Figure 26: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers for examples with $\{\neq\mathcal{L},\neq r,=s\}$. Top is XGLM, bottom mT5.
+F.3 Different Relation, Same Subject
+Figure 40: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers for examples with $\{\neq\mathcal{L},\neq r,=s\}$. Models from top to bottom: XGLM, EuroLLM, mT5.
 Figure 41: Probability of the $\mathcal{L}_{p}(o_{p})$ when patching at different layers in XGLM, for examples with $\{\neq\mathcal{L},\neq r,=s\}$.
 Figure 42: Probability of the $\mathcal{L}_{c}(o_{c})$ when patching at different layers in XGLM, for examples with $\{\neq\mathcal{L},\neq r,=s\}$. Note that the plots do not share the y-axis.
+Figure 43: Probability of the $\mathcal{L}_{p}(o_{p})$ when patching at different layers in EuroLLM, , for examples with $\{\neq\mathcal{L},\neq r,=s\}$.
+Figure 44: Probability of the $\mathcal{L}_{c}(o_{c})$ when patching at different layers in EuroLLM, for examples with $\{\neq\mathcal{L},\neq r,=s\}$. Note that the plots do not share the y-axis.
 Figure 45: Probability of the $\mathcal{L}_{p}(o_{p})$ when patching at different layers in mT5, , for examples with $\{\neq\mathcal{L},\neq r,=s\}$.
 Figure 46: Probability of the $\mathcal{L}_{c}(o_{c})$ when patching at different layers in mT5, for examples with $\{\neq\mathcal{L},\neq r,=s\}$. Note that the plots do not share the y-axis.
-Figure 31: Number of times the patch object is predicted in the context language for the experiment of same relation different subject. Left XGLM, right mT5.
-Figure 32: Number of times the object is predicted in the opposite language in mT5 in the different relation same subject experiment.
-Figure 33: Number of times the context object is predicted in the patch language in XGLM in the experiment of different relation same subject.
+Figure 47: Number of times the patch object is predicted in the context language for the experiment of same relation different subject. Models from top to bottom: XGLM, EuroLLM, mT5.
+Figure 48: Number of times the object is predicted in the opposite language in mT5 in the $\{\neq r,=s,\neq\mathcal{L}\}$ experiment.
+Figure 49: Number of times the context object is predicted in the patch language in the $\{\neq r,=s,\neq\mathcal{L}\}$ experiment. Left XGLM, right EuroLLM.

### v3

@@
 These insights reveal unique mechanisms in multilingual LLMs for recalling information, highlighting the need for new methodologies—such as knowledge evaluation, fact editing, and knowledge acquisition—that are specifically tailored for multilingual LLMs.
+[^1] https://github.com/constanzafierro/multilingual_factual_memorization
 How Do Multilingual Language Models Remember Facts?
-Constanza Fierro∗†   Negar Foroutan   Desmond Elliott   Anders Søgaard University of Copenhagen    EPFL
-**footnotetext: Correspondance: Constanza Fierro <c.fierro@di.ku.dk>.
+Constanza Fierro   Negar Foroutan   Desmond Elliott   Anders Søgaard Department of Computer Science, University of Copenhagen EPFL
+**footnotetext: Correspondence: Constanza Fierro <c.fierro@di.ku.dk>.
 1 Introduction
@@
 Such assessments will be critical for determining the trustworthiness of multilingual LLMs, if trust relies on knowledge (Grasswick, 2010; Hawley, 2012; Nguyen, 2022).
+Figure 1: Aggregated patching results (§6) for EuroLLM.
+In propositional factual statements, e.g.
+“Paris is the capital of”, the last token representation contains the function that solves the task.
+This function is formed in two stages: first, the relation to extract is encoded (green), and then the language is composed into the function (purple).
+Finally, the function is applied to the subject, and the predicted object is resolved (blue).
 Mechanistic interpretability research has begun uncovering how models store and retrieve knowledge internally, with recent studies identifying specific components for knowledge storage (Meng et al., 2022; Sharma et al., 2024) and retrieval mechanisms (Geva et al., 2023; Chughtai et al., 2024).
@@
 (3) When does language play a role in the recall mechanism?
-To address the first question, we use causal tracing analysis to assess if early MLP modules processing the final subject token are as decisive in multilingual models as in English-centric models (Meng et al., 2022).
-Our results (§3) show that EuroLLM and mT5 exhibit strong causal effects for all subject tokens, with EuroLLM exhibiting this in earlier layers and mT5 across all encoder layers.
-In XGLM, the last subject token has a stronger effect, but with two early sites, unlike English GPT’s single site.
-Additionally, in all models, MLPs in later layers are decisive in recovering factual information, a contrast to Meng et al.
-(2022) findings.
-Next, we investigate the second question by analyzing the mechanisms described in Geva et al.
-(2023); namely, the three-step process, where the relation information flows to the last token, then the subject flows, and finally the attention layers extract the object (§4).
-We find that in multilingual decoder-only models the information flows similarly to monolingual models, but differently for the encoder-decoder model mT5.
-In terms of the extraction event, in decoder-only LLMs, both feed-forward and attention sublayers contribute to it, unlike English autoregressive models, where attention modules dominate.
-In mT5, this mechanism is primarily performed by the cross-attention.
+To address the first question, we use causal tracing analysis to assess if early MLP modules processing the final subject token are as decisive in multilingual models as in English-centric ones (Meng et al., 2022).
+Our results (§4) show that EuroLLM and mT5 exhibit strong causal effects for the last subject token—EuroLLM in earlier layers, and mT5 across all encoder layers.
+Additionally, in all models, both MLPs and MHSAs in later layers play a decisive role in recovering factual information—unlike in Meng et al.
+(2022), where only MHSAs were active at the late site.
+Next, we investigate the second question by analyzing the three-step process described by Geva et al.
+(2023): the relation information flows to the final token, followed by subject information, and finally, the attention layers extract the object (§5).
+We find that in multilingual LLMs, subject information flows similarly to monolingual ones, but non-subject token flow differs, and this analysis alone cannot determine the path of relation information.
+Regarding the final extraction, both feed-forward and attention sublayers contribute in decoder-only LLMs—unlike in English autoregressive LLMs, where attention modules dominate.
+In mT5, this mechanism is primarily handled by cross-attention.
 Overall, our findings indicate that some of the localization (§4) and mechanisms (§5) of fact retrieval in English LLMs generalize to multilingual LLMs, but with key variations.
-Finally, to address the third question, we investigate where language plays a role within the three-step process described by Geva et al.
-(2023), to characterize how and where factual knowledge cross-lingual transfer may occur.
-Using activation patching (Zhang and Nanda, 2024; Ghandeharioun et al., 2024) we insert the intermediate representation of the last token from an English forward pass into the forward pass of another language (§5).
+Finally, to address the third question, we investigate where language plays a role within the three-step process, to characterize how and where factual knowledge cross-lingual transfer may occur.
+Using activation patching (Zhang and Nanda, 2024) we insert the intermediate representation of the last token from an English forward pass into the forward pass of another language (§6).
 Our results reveal that the last token acts as a Function Vector (FV) Todd et al.
@@
 Crucially, the fact that the FV formed in one language can be used with an input in another language demonstrates that the subject and relation representations are largely language-independent, while the extraction event is language-specific.
-Furthermore, in decoder-only LLMs (XGLM, EuroLLM), the FV is constructed in two distinct phases: first, it encodes only the relation, and later, the language is incorporated.
+Furthermore, in decoder-only LLMs, the FV is constructed in two distinct phases: first, it encodes only the relation, and later, the language is incorporated (Figure 1).
 These findings advance our understanding of factual recall, with implications for cross-lingual knowledge transfer, knowledge editing, and trustworthiness in multilingual LLMs.
@@
 Additionally, the late-site causal effect of MLPs and their dominance in the extraction phase suggest that fact editing techniques and knowledge evaluations must extend beyond early MLPs and attention layers (Tamayo et al., 2024).
+2 Related Work
+Factual Knowledge Recall
+Petroni et al.
+(2019) analyzed factual knowledge in pre-trained language models using the LAMA dataset, which pairs cloze-test templates with WikiData triplets.
+[^3] One might object to treating cloze-test performance as indicative of knowledge, given LLM inconsistencies.
+However, following Fierro et al.
+(2024), we adopt this view, as LLMs are often consistent and can sometimes justify responses through world models or training data attribution.
+This was extended to multilingual settings by translating LAMA (Kassner et al., 2021; Jiang et al., 2020).
+Later, Elazar et al.
+(2021) evaluated the consistency of such knowledge by manually curating paraphrases of LAMA to create ParaRel.
+Further work examined factual consistency cross-lingually (Fierro and Søgaard, 2022; Qi et al., 2023), using machine-translated ParaRel templates and WikiData-based subject/object translations to produce mParaRel.
+Interpretability on Factual Knowledge Recall
+Early studies by Meng et al.
+(2022) and Geva et al.
+(2023) mechanistically analyzed knowledge localization and information flow in GPT models (Brown et al., 2020), using English data.
+Later, Chughtai et al.
+(2024) showed that answers emerge from summing independent components in GPT and Pythia models, while Sharma et al.
+(2024) extended these findings to Mamba (Gu and Dao, 2023), also focusing on English.
+Activation Patching
+Ghandeharioun et al.
+(2024) introduced Patchscope, a framework for decoding intermediate representations by patching them into forward passes.
+Dumas et al.
+(2024) applied this method to last-token representations in a translation task, showing that models encode language-agnostic concepts—consistent with our findings and those of Foroutan et al.
+(2022).
+Wang et al.
+(2024) similarly found that the last token encodes the relation in factual English queries at a specific stage of computation.
+We extend these findings to a cross-lingual factual recall setting.
+While Dumas et al.
+(2024) concluded that models first resolve the output language, we find that models first resolve the query’s relation, then the language.
+We thus interpret the last token as a function vector (Todd et al., 2024), which encodes the task-specific function—e.g., translating in the case of Dumas et al.
+(2024).
+Additionally, we show that in decoder-only models, the FV composes the output language onto the relation function vector.
 3 Experimental Setup
 We focus on a simple form of factual knowledge recall, where LMs are tasked with predicting the correct object $o$ for a given subject $s$ and a relation $r$.
-These $(s,r,o)$ triplets are obtained from WikiData, and natural language templates are used to describe the relation, with placeholders for the subject and object.222 For example, the relation born-in could use the template “[X] was born in [Y]”, where [X] is the subject and [Y] is the object to be predicted.
+These $(s,r,o)$ triplets are obtained from WikiData, and natural language templates are used to describe the relation, with placeholders for the subject and object.444 For example, the relation born-in could use the template “[X] was born in [Y]”, where [X] is the subject and [Y] is the object to be predicted.
 For our analysis, we select 10 typologically diverse languages representing various scripts, families, and word orders: English (en), Spanish (es), Vietnamese (vi), Turkish (tr), Russian (ru), Ukrainian (uk), Japanese (ja), Korean (ko), Hebrew (he), Persian (fa), and Arabic (ar).
@@
 [^5] For decoder-only models, we only use the templates in mParaRel that have the object placeholder at the end of the sentence, while for encoder-decoder, we use all the templates.
-The decoder-only LLMs are XGLM (Lin et al., 2021),
+The decoder-only LLMs are XGLM (Lin et al., 2021), with 7.5B parameters and 32 layers, and EuroLLM (Martins et al., 2024) with 9B parameters and 42 layers; the encoder-decoder mT5-xl (Xue et al., 2021) has with 3.7B parameters and 24 encoder-decoder layers.
+The pre-training data varies across models: mT5 is pretrained on 101 languages, covering all the languages in our study; XGLM covers 30 languages, excluding uk, he, and fa; while EuroLLM covers 35 languages, excluding vi, he, and fa.
+Data
+We use the mParaRel dataset (Fierro and Søgaard, 2022), which includes triplets and templates for 45 languages.
+[^6] We augment the objects in mParaRel, filter out trivial examples, and when there are enough examples, we use a crosslingual subset (see Appendix A).
+Figure 2: Average indirect effect (IE) on logit scores, when the subject input is corrupted and some activations are restored to their uncorrupted values.
+The $y$ axis indicates the layer where the restoration was performed, and the $x$ axis specifies the token(s) over which we average the IE.
+The MHSA and MLP are restored in windows of 12% consecutive layers.
+Top GPT2-XL (only English data), bottom EuroLLM (XGLM and mT5 in Figure 11).
+To investigate the process of knowledge recall we only consider examples where the model predicts the correct object completion (Meng et al., 2022; Geva et al., 2023).
+Since mParaRel provides multiple paraphrased templates for each relation, we greedy-decode for every available template corresponding to a given triplet and check for an exact match.
 XGLM EuroLLM mT5
@@
 Table 1: Number of facts $(s,r,o)$ correctly predicted.
-with 7.5B parameters and 32 layers, and EuroLLM (Martins et al., 2024) with 9B parameters and 42 layers; the encoder-decoder mT5-xl (Xue et al., 2021) has with 3.7B parameters and 24 encoder-decoder layers.
-The pre-training data varies across models: mT5 is pretrained on 101 languages, covering all the languages in our study; XGLM covers 30 languages, excluding uk, he, and fa; while EuroLLM covers 35 languages, excluding vi, he, and fa.
-Data
-We use the mParaRel dataset (Fierro and Søgaard, 2022), which includes triplets and templates for 45 languages.
-These templates are machine translations of the ParaRel English templates (Elazar et al., 2021).
-[^4] We augment the objects in mParaRel, filter out trivial examples, and when there are enough examples, we use a crosslingual subset (see Appendix B).
-To investigate the process of knowledge recall we only consider examples where the model predicts the correct object completion (Meng et al., 2022; Geva et al., 2023).
-Since mParaRel provides multiple paraphrased templates for each relation, we greedy-decode for every available template corresponding to a given triplet and check for an exact match.
 If multiple templates yield a match, we randomly select one for the analysis.
-In cases where an article or other filler tokens precede the object, we include these tokens in the input text to ensure that when the example is fed into the model for our analysis, the next predicted token is the first token of the object (Implementation details in Appendix B.1).
+In cases where an article or other filler tokens precede the object, we include these tokens in the input text to ensure that when the example is fed into the model for our analysis, the next predicted token is the first token of the object (Implementation details in Appendix A.1).
 Table 1 presents the number of examples for which the correct object is predicted.
@@
 (2022) and add $\epsilon\sim\mathcal{N}(0,(3\sigma)^{2})$, with $\sigma$ being the standard deviation of the subjects tokens embeddings from the data used.
+We repeat the experiment ten times with different noise samples, and report the average.
 and observe the corrupted probability ~$\tilde{\mathbb{P}}(o)$ of the originally predicted token.
@@
 In other words, multiple layers contribute the same behavior to the residual, and their sum produces the observed effect.
-We set $w$=$4$ for XGLM, $w$=$5$ for EuroLLM, and $w$=$3$ for mT5, to restore $\sim 12\%$ layers.
-We repeat the corrupted run with restoration ten times with different noise samples, and report the average.
+We restore windows of $12\%$ consecutive layers, so $w$=$4$ for XGLM, $w$=$5$ for EuroLLM, and $w$=$3$ for mT5.
 We compare our results in multilingual LLMs to those of Meng et al.
-(2022), who analyzed the factual recall of GPT-2 XL in English [^7] They used the CounterFact dataset, which is based on ParaRel.
-Thus, their English data is the same as our data.
-and reached two key conclusions: (1) they identified an “early site”, where the MLPs processing the last subject token in the early and middle layers play a crucial role in recovering from input corruption; and (2), they found a “late site”, where the attention modules processing the last token in the later layers also significantly contribute to prediction recovery.
+(2022), who analyzed the factual recall of GPT-2 XL in English and reached two key conclusions: (1) they identified an “early site”, where the MLPs processing the last subject token in the early and middle layers play a crucial role in recovering from input corruption; and (2), they found a “late site”, where the attention modules processing the last token in the later layers also significantly contribute to prediction recovery.
 [^9] While they consider the late site unsurprising, as it directly precedes the final prediction, this observation primarily applies to hidden state restoration, not necessarily attention restoration.
@@
 Each layer represents the effect of the knockout on a window of $w$ layers.
-Top EuroLLM ($w=7$), bottom mT5 ($w=4$).
-XGLM in Figure 21.
-We present the causal analysis plots for each model and language in Appendix C and present our main observations here.
+Top is the average, bottom EuroLLM ($w=7$).
+XGLM and mT5 in Figure 24.
+We present the average causal analysis results in Figure 2 (results per language in Appendix B).
 [^10] In line with Zhang and Nanda (2024) we find that analyzing the $\text{IE}_{\mathbb{P}}$ overestimates the causal effect of some tokens over others.
-For example, in XGLM by looking at the $\text{IE}_{\mathbb{P}}$ we would not find an early site, which is present when studying the $\text{IE}_{LS}$ (see Figure 11 vs Figure 9), or in EuroLLM the last subject token would seem more relevant than the other subject tokens (see Figure 15 vs Figure 13).
+For example, in EuroLLM the last subject token MHSA seem more relevant than the last token MLPs (see Figure 18 vs Figure 16).
 So we base our main observations using $\text{IE}_{LS}$.
-Our results indicate that the last subject token does not consistently have a stronger effect on information recovery in early MLPs.
-In EuroLLM and mT5, all subject tokens exhibit similar effects, while in XGLM, the last subject token has a stronger influence.
-In EuroLLM, the early site spans half of the initial layers, whereas in mT5, it covers all encoder layers.
-For XGLM, the early site occurs twice: once in the initial layers and again in the middle layers.
-Regarding the late site, we observe a strong causal effect in MLP layers before the very last layers in all models.
-The layers where these sites occur are consistent across languages, with only a slight variation in the late site’s endpoint for EuroLLM.
-These findings suggest that some, but not all, conclusions from English LLMs generalize to multilingual LLMs.
-The MLP early site extends across models, and the MLP late site, which was not observed by Meng et al.
-(2022), was present in all models studied.
-This has implications for knowledge localization and fact editing methodologies, which should account for both early and late MLPs in multilingual contexts.
+Our results indicate that the early site in MLPs—centered on the last subject token—is also present in multilingual LLMs.
+We observe this causal effect in both EuroLLM and mT5.
+In mT5, the first subject token also shows causal relevance, though marginally less so, and the early site spans all encoder layers.
+Across languages, the patterns are highly consistent.
+Interestingly, we do not observe an MLP early site in XGLM, either for English or in the MHSAs.
+As for the late site, we find a strong causal effect in both MLP and MHSA layers just before the final layers in all models.
+The position of the late site is consistent across languages, with only slight variation in its endpoint for EuroLLM.
+These findings suggest that some conclusions from English LLMs generalize to multilingual ones, but not all.
+The early MLP site is consistent across models, while at the late site, both MLPs and MHSA are important in multilingual LLMs—unlike in English LLMs, where MHSA dominates.
+This has implications for knowledge localization and fact editing, which should consider both early and late MLPs in multilingual contexts.
 5 Factual Recall Components
@@
 Information Flow
-[…]
-Figure 1 presents the results with plots per language in Appendix D.
-Firstly, the results show that in each model the information flows fairly similarly for all the languages, and similar patterns to those found in English GPT are present (Geva et al., 2023).
+We use attention knockout (Geva et al., 2023) to analyze how information propagates to the final token.
+This method involves intervening in the attention mechanism: for XGLM and EuroLLM, we modify the last token’s self-attention, while for mT5, we intervene in the decoder’s cross-attention for the final token.
+We knock out attention connections by setting the attention scores to zero between the final token and a token set $\{t\}$, which can be: subject tokens, non-subject tokens, or the last token itself.
+Following Geva et al.
+(2023) we apply this intervention across consecutive layers within a window of size $w$ centered on layer $l$, covering $18\%$ of the total layers ($w$=$6$ for XGLM, $w$=$7$ for EuroLLM, and $w$=$4$ for mT5).
+The relative probability change is given by (~$(\tilde{\mathbb{P}}(o)-\mathbb{P}(o))/\mathbb{P}(o)$, where ~$\tilde{\mathbb{P}}(o)$ is the probability of the originally predicted token $o$ after attention knockout .A significant drop indicates that the knocked-out tokens contribute critically to the final token’s prediction at that layer.
+In Figure 3 we present average results, plots per language can be found in Appendix C.
+The results show that in each model the information flows fairly similarly for all the languages, and somewhat similar patterns to those in English GPT2-xl.
 On the one hand, the subject information flows to the last token most critically at the later layers in all models.
-On the other hand, the relation flows throughout all the layers, with EuroLLM and mT5 having more similar curves.
-Nevertheless, we note two differences from the conclusions reached with English GPT (Geva et al., 2023): (1) the propagation of non-subject tokens to the last token does not strictly precedes the subject propagation, in XGLM the drop is the highest in the earlier layers but it continues to be important until the end, and in mT5 and EuroLLM this information has two peaks, in the early and later layers; and (2) the last token of mT5 encodes critical information that flows from one layer to the next through the cross-attention (as opposed to the negligible flow of the last token in the decoder-only models).
+On the other hand, the non-subject tokens information flows throughout all the layers, with EuroLLM and mT5 having more similar curves.
+Nevertheless, we note two differences from the conclusions reached with English GPT (Geva et al., 2023): (1) the propagation of non-subject tokens to the last token does not strictly precedes the subject propagation, in XGLM the drop is the highest in the earlier layers but it continues to be important until the end, and in mT5 and EuroLLM this information has two peaks, in the early and later layers; and (2) the last token of mT5 encodes critical information that flows from one layer to the next through the cross-attention (as opposed to the negligible flow of the last token in decoder-only models).
 [^11] We hypothesize this may occur because the last token encodes which sentinel token is being generated, indicating where to fill in the input.
 Since we see an almost identical curve when the last token cannot attend to the sentinel token in the decoder (Figure 27).
-Figure 2: Extraction rates. Left XGLM, right EuroLLM, bottom mT5.
+Figure 4: Extraction rates. Fraction of examples where the top-1 token under the vocabulary projection matches the final output token.
 Prediction Extraction
@@
 The extraction rate is the proportion of examples for which an extraction event occurs at a given layer.
-Figure 2 shows the extraction rates for each model, with a detailed language breakdown in Appendix E.
-Our results demonstrate that extraction events can be detected in multilingual LLMs, though rates vary across languages.
+Our results demonstrate that extraction events can be detected in multilingual LLMs (Figure 4), though rates vary across languages (Appendix D).
+[^12] This variance may result from the strict criteria applied during the extraction event analyses, where only top-1 matches were considered.
 A key finding is the prominence of MLP modules in object extraction for multilingual decoder-only models (XGLM and EuroLLM).
 To rule out the possibility that MLPs simply forward extracted objects from preceding attention layers, we measure how often MLPs perform an extraction without prior attention extraction (Figure 31).
-We find that MLPs indeed perform the extraction for most languages, though in English, attention modules dominate, aligning with GPT-2 results (Geva et al., 2023).
-In mT5, cross-attention layers drive the extraction, with MLPs playing a secondary role, resembling GPT-2’s behavior.
+We find that MLPs indeed perform the extraction for most languages, though in English, attention modules dominate, aligning with GPT2-xl results.
+In mT5, cross-attention layers drive the extraction, with MLPs playing a secondary role, resembling GPT’s behavior.
 Additionally, in mT5 the extraction events occur in later layers, with peaks in 19 and 22.
-The results show that in multilingual LLMs, information flow from non-subject tokens to the final token is less localized compared to English LLMs, possibly because the model processes relational and language information at different stages.
-Additionally, the extraction mechanism in multilingual LLMs is more complex, involving both attention and MLP modules.
+The results show that in multilingual LLMs the extraction mechanism is more complex, involving both attention and MLP modules.
 This implies that editing techniques should focus not only on early MLPs enhancing subject representations but also on the later MLPs that extract the object.
+Figure 5: Patching strategy. The patch example’s last token is inserted in the context example’s forward pass, which perturbs the output of subsequent layers.
 6 Language Information Flow
@@
 Here, we interpret the last token representation as an FV because it triggers a specific execution for different contexts..
-Figure 3: Patching strategy. The patch example’s last token is inserted in the context example’s forward pass, which perturbs the output of subsequent layers.
-(a) Counts of examples where the patch produces the object prediction.
-(b) Histogram of patches that cause the context answer to be predicted in the patch language, $\mathcal{L}_{p}(o_{c})$.
-Figure 5: Prediction tokens from (a) $\{=\mathcal{L},\neq r,\neq s\}$ and (b) $\{\neq\mathcal{L},=r,\neq s\}$ for XGLM, EuroLLM, mT5.
+(a) Number of examples where the patch produces one of the four valid objects.
+The green curve (middle) represents layers where the last token’s representation encodes the relation but not yet the object.
+(b) Number of examples where the patch causes the output to be $\mathcal{L}_{p}(o_{c})$ (the context object in the patch language).
+The last token’s representation at these layers encodes the relation to extract and in which language, but not yet the object.
+Figure 7: Patches per layer triggering a specific object prediction. (a) Setup $\{=\mathcal{L},\neq r,\neq s\}$; (b) $\{\neq\mathcal{L},=r,\neq s\}$.
 Let $h^{l}_{t}$ be the representation at layer $l$ of the last token in the input.
 For a given input $p$, we take $h^{l}_{t,p}$ and patch it into the forward pass of another input $c$ at the same layer; that is, we set $h^{l}_{t,c}\leftarrow h^{l}_{t,p}$, and then continue the forward pass.
-We refer to the example $p$ as the patch example, and $c$ as the context example (see Figure 3).
-By patching the last token representation in each layer and measuring the model’s predicted output, we can study when the representation contains the information about relation, language, and predicted object, and how cross-lingual these representations are.
+We refer to the example $p$ as the patch example, and $c$ as the context example ( Figure 5).
+By patching the last token representation in each layer and observing the model’s predicted output, we can study when the representation contains the information about relation, language, and predicted object, and how cross-lingual these representations are.
+In all experiments, English is used as the patch language.
 An input example expresses a relation $r$ and a subject $s$ in language $\mathcal{L}$, which we note $(\mathcal{L}(r),\mathcal{L}(s))$.
-We conduct three experiments where the patch and context examples share certain input characteristics.
-In particular, we test: (1) $\{=\mathcal{L},\neq r,\neq s\}$, where the language is the same but the relation and subject differ; (2) $\{\neq\mathcal{L},=r,\neq s\}$, where the relation remains the same but the language and subject differ; and (3) $\{\neq\mathcal{L},\neq r,=s\}$, where the subject is the same but the language and relation differ.
+We conduct three experiments where the $p$ and $c$ examples share certain input characteristics: (1) $\{=\mathcal{L},\neq r,\neq s\}$, the language is the same but the relation and subject differ; (2) $\{\neq\mathcal{L},=r,\neq s\}$, the relation remains the same but the language and subject differ; and (3) $\{\neq\mathcal{L},\neq r,=s\}$, the subject is the same but the language and relation differ.
 (1) allows us to study how the relation is encoded while controlling for the language, whereas (2)-(3) help us explore how the interaction between relation and language affects the extraction of the object from the subject.
-In all experiments, English is used as the patch language, $\mathcal{L}_{p}=\text{en}$.
+In Figure 1 and 35 we show the aggregated results of these 3 settings, with setup (1) we localize when the object is resolved (blue), with setup (2) when the relation and language are encoded (purple), and with setup (3) when the relation is encoded (green).
 Let the patch input be $(\mathcal{L}_{p}(r_{p}),\mathcal{L}_{p}(s_{p}))$ and the context input be $(\mathcal{L}_{c}(r_{c}),\mathcal{L}_{c}(s_{c}))$.
@@
 For predicted tokens, we check if they match the patch or context object, or a variant with swapped language or relation (e.g., if $\mathcal{L}_{p}(o_{c})$ is predicted in the setups (2)-(3)).
-We only consider valid patch-context pairs where $o_{r_{p},s_{c}}$ and $o_{r_{c},s_{p}}$ exist, and for predicted token analysis, we enforce distinct spellings.
-[^13] E.g.
+[^15] We only consider valid patch-context pairs where $o_{r_{p},s_{c}}$ and $o_{r_{c},s_{p}}$ exist, and for predicted token analysis, we enforce distinct spellings, e.g.
 to claim the predicted token is $t=\mathcal{L}_{p}(o_{c})$ we require $t\neq\mathcal{L}_{c}(o_{c})$, as languages may share spellings (e.g., “Asia” in English and Spanish).
@@
 See Table 4 and 6 for the number of examples for each model and experiment.
-Figure 6: Relative probability change of $\mathcal{L}_{c}(o_{c})$ when patching $\{\neq\mathcal{L},=r,\neq s\}$, for XGLM, EuroLLM, mT5.
-Figure 7: Relative probability change of $\mathcal{L}_{c}(o_{c})$ when patching $\{\neq\mathcal{L},\neq r,=s\}$, for XGLM, EuroLLM, mT5.
+Figure 8: Percentage change in the probability of the context object $\mathbb{P}(\mathcal{L}_{c}(o_{c}))$ when patching $\{\neq\mathcal{L},=r,\neq s\}$.
+Figure 9: Percentage change in the probability of the context object $\mathbb{P}(\mathcal{L}_{c}(o_{c}))$ when patching $\{\neq\mathcal{L},\neq r,=s\}$.
 Different Relation, Different Subject
@@
 We sample 1000 examples for which mParaRel has the objects $o_{r_{c},s_{p}}$ and $o_{r_{p},s_{c}}$.
-We present the prediction results in Figure 5(a), and probability plots in Figure 32.
-The relation information is contained in the last token representation when the green curve increases, since patching at that point starts producing the object corresponding to the relation in the patch $o_{r_{p},s_{c}}$.
-When the green curve decreases and the blue curve increases, the object has been fully extracted and is encoded in the last token vector, as predictions from these layers correspond to $o_{r_{p},s_{p}}$ and no information is taken from the context $c$.
-The layers where this happens align with the peaks of the extraction rate in English (Figure 25-27).
-This shows that the extraction measured by the vocabulary projection represents the point at which the object is extracted and encoded in the last token representation.
-If the object was encoded earlier and only decoded with the extraction event,$o_{r_{p},s_{p}}$ would be predicted from patches at earlier layers.
+We present the prediction results in Figure 7(a) and the probability plots in Figure 36.
+An increase in the green curve indicates that relation information becomes available in the last token representation, as patching at that layer produces $o_{r_{p},s_{c}}$—the object corresponding to the relation in the patch example.
+When the green curve declines and the blue curve rises, the object has been fully extracted and encoded in the last token, since predictions now yield $o_{r_{p},s_{p}}$ and are no longer influenced by $c$.
+These transitions align with peaks in the extraction rate for English (Figure 28-30).
+This confirms that extraction, as measured by the vocabulary projection, marks the point where the object is both available and encoded in the last token representation.
+If the object had been encoded earlier and only decoded at the extraction point, $o_{r_{p},s_{p}}$ would have been predicted from earlier-layer patches.
 Same Relation, Different Subject
@@
 On the other hand, if we obtain $\mathcal{L}_{c}(o_{p})$ (“París”), we could infer that the language is encoded after the extraction.
-We present the probability of the context answer token $\mathcal{L}_{c}(o_{c})$ in Figure 7.141414The probability of $\mathcal{L}_{p}(o_{p})$ and per language plots are provided in Appendix F.2).
+We present the probability of the context answer token $\mathcal{L}_{c}(o_{c})$ in Figure 9.161616The probability of $\mathcal{L}_{p}(o_{p})$ and per language plots are provided in Appendix E.2.
 We observe that for decoder-only models, patching in the early layers generally hurts the model’s performance for most languages, while, patching in the middle layers either increases the probability or results in only a minimal decrease.
@@
 In the case of mT5, however, the probability of $\mathcal{L}_{c}(o_{c})$ consistently decreases, plateauing in the middle layers before dropping to zero.
-From the attention knockout analysis, we recall that in mT5, the subject is integrated into the last token representation only after layer 15, while the relation is consistently represented throughout the middle layers.
+From the attention knockout analysis, recall that in mT5 the subject is integrated into the last token representation only after layer 15, while the relation is consistently represented throughout the middle layers.
 Therefore, these patching results imply that the relation encoded in the last token from the patch input does not help in retrieving the correct context object in the context language in mT5, whereas it proves useful in XGLM and EuroLLM.
@@
 To analyze if this is the case, we now apply patching on examples with different languages and relations but the same subject $\{\neq\mathcal{L},\neq r,=s\}$, e.g., “France’s capital is” and “El idioma oficial de Francia es” (Gloss: “The official language of France is”).
-We observe that the probability of $\mathcal{L}_{c}(o_{c})$ (Figures  7) in decoder-only models, unlike the previous experiment, decreases early for all languages (except tr and ko), plateaus around the middle layers, and then drops to zero by the mid-layer range.
+We observe that the probability of $\mathcal{L}_{c}(o_{c})$ (Figures  9) in decoder-only models, unlike the previous experiment, decreases early for all languages (except tr and ko), plateaus around the middle layers, and then drops to zero by the mid-layer range.
 For mT5, the probability drops at the beginning but, instead of plateauing as before, it continues to decline until it reaches zero.
@@
 More detailed analysis of what is being attended when the relation flows and when the language flows should be performed in future work.
-Figure 8: Patches that cause the prediction to be $\mathcal{L}_{c}(o_{p})$. Top XGLM, bottom EuroLLM.
+Figure 10: Patches that cause the prediction to be $\mathcal{L}_{c}(o_{p})$, which shows the layers where the last token representation contains the relation but not yet the language.
 7 Conclusion
-In this paper, we analyzed factual knowledge recall mechanisms in 10 languages using multilingual transformer-based LMs, comparing them to prior research on English recall in monolingual autoregressive LLMs.
+In this paper, we analyzed factual knowledge recall mechanisms in 10 languages using multilingual transformer-based LMs, comparing them to recall in English autoregressive LLMs.
 We discovered that some mechanisms, such as the flow of subject representations in the later layers and the extraction phase, are present in multilingual and monolingual models.
-However, we also identified notable differences, including the localization of knowledge in not only one but two MLP sites; and the joint role of MLPs and attention modules during the extraction phase in multilingual models.
+However, we also identified notable differences, including the join role of late MLPs and attention modules during the extraction phase in multilingual models.
 A key contribution of our work is the first-ever investigation of language encoding during recall, achieved through patching representations.
@@
 Contributing with new directions for the study of cross-lingual transfer and knowledge localization.
-7 Limitations
+Limitations
 In this paper, we examined three model architectures, leaving out the effects of model sizes, instruction fine-tuning, or models like Llama that can behave multilingually but have less coverage and less multilingual pre-training data.
@@
 Lastly, we described the main mechanisms found in XGLM, EuroLLM and mT5, however other weaker mechanisms could be at play, which could describe, for example, the low extraction rates found for some languages (Figure 30) or the few examples where the object seems to be encoded from early layers before the extraction takes place (Figure 51).
-Appendix A Related Work
-Factual Knowledge Recall
-Petroni et al.
-(2019) analyzed factual knowledge in pre-trained language models, by creating the LAMA dataset, which contains natural language templates and WikiData triplets.
-Subsequently, Elazar et al.
-(2021) developed the ParaRel dataset, which contains manually paraphrases of the LAMA templates, to measure the consistency of factual knowledge.
-[^15] One may object to calling the ability to solve cloze tests as knowledge.
-LLMs are sometimes inconsistent and cannot always justify the propositions they generate.
-We side with Fierro et al.
-(2024) in thinking knowledge may nevertheless be the appropriate concept, since LLMs are mostly consistent and can sometimes provide justification, e.g., in virtue of world models or training data attribution.
-This led to further research in multilingual settings, Kassner et al.
-(2021); Jiang et al.
-(2020) translated LAMA, creating mLAMA and X-FACTR, and evaluated the factual knowledge acquired by multilingual pre-trained models, while Wang et al.
-(2024a) evaluated knowledge from the perspective of knowledge edits and their impact between Chinese and English (Zhu et al., 2021).
-Building on this, Yin et al.
-(2022) curated a dataset that focuses on the cultural aspect of factual knowledge.
-While Fierro and Søgaard (2022) studied multilingual consistency of knowledge using mParaRel, a machine-translated version of the English templates in ParaRel, where subjects and objects are translated using WikiData identifiers.
-In a related effort, Qi et al.
-(2023) measured the cross-lingual consistency of multilingual LLMs by evaluating the consistency across languages, rather than just across paraphrases in the same language.
-Qi et al.
-(2023) used the mLAMA dataset.
-In this paper, we used mParaRel since it is an extended version of mLAMA, as it is a translation of the paraphrases in mParaRel.
-Interpretability on Factual Knowledge Recall
-As discussed earlier, Meng et al.
-(2022) was among the first to analyze, from a mechanistic perspective, whether a model’s knowledge could be localized; while Geva et al.
-(2023) conducted the first study on how information flows within a model to construct the final predicted object representation.
-Both studies focused on the GPT models (Brown et al., 2020), using English-language data.
-Later, Chughtai et al.
-(2024) contributed by demonstrating that models arrive at an answer through several independent components, with the sum of these components yielding the correct object.
-Their analysis covered both GPT and Pythia models, also using English data.
-Subsequently, Sharma et al.
-(2024) applied the mechanisms identified by Meng et al.
-(2022) and Geva et al.
-(2023) to the state-space language model Mamba (Gu and Dao, 2023), focusing on facts expressed in English.
-Additionally, Hernandez et al.
-(2024) examined whether a linear transformation of the subject is sufficient to represent various relations, analyzing GPT and Llama models with English data.
-Activation Patching
-Ghandeharioun et al.
-(2024) introduced the Patchscope framework to study what an intermediate representation encodes.
-In their experiments, they focused on “decoding” the contents of a latent representation by patching it into the forward pass of another input.
-More closely related to our work, Dumas et al.
-(2024) patched the last token representation to investigate if models encode language-agnostic concepts in a word translation task.
-Their findings, consistent with ours and those of Foroutan et al.
-(2022), support the existence of language-agnostic representations.
-Similarly, Wang et al.
-(2024b) patched the last token representation to study how relations are encoded in factual English queries, finding that the last token encodes the query’s relation at a specific computational stage—aligning with our results.
-In this paper, we are able to draw more general conclusions given that we study a cross-lingual factual recall setup.
-While Dumas et al.
-(2024) concluded that LLMs first resolve the output language, we find that the model first resolves the query’s relation and then the language.
-Thus, we interpret the last token representation as a Function Vector (FV) (Todd et al., 2024), encoding the function to solve the given task: for Dumas et al.
-(2024) the task was translation so the function encodes the language to which translate.
-Finally, we are also able to show that for decoder-only models the FV composes the language onto the relation FV, which extends the findings from Wang et al.
-(2024b).
-Multilingual Language Models
-In this work, we focused on generative models and selected mT5 (Xue et al., 2021), XGLM Lin et al.
-(2021), and EuroLLM (Martins et al., 2024) due to their broad language coverage, highly multilingual training data, and pre-training with up-sampling of low-resource languages to ensure more balanced language representation.
-However, many other models are available.
-For instance, BLOOM (Le Scao et al., 2023) is a decoder-only model trained on 46 languages, although low factual recall accuracy has been reported (Qi et al., 2023).
-Llama-2 (Touvron et al., 2023) has been used in some multilingual analyses (Wendler et al., 2024; Dumas et al., 2024), but it is primarily an English-centric model given that 90% of its pre-training data is in English.
-More recent models that could have been considered include mGPT (Shliazhko et al., 2022), a decoder-only model trained on 61.
-On the other hand, the Aya-101 model (Üstün et al., 2024) is an instruction fine-tuned version of mT5; however, we focus on pre-trained multilingual LMs to compare our analysis to the English monolingual pre-trained versions previously studied.
-Language Script Family SOV/SVO
-English Latin Germanic SVO
-Spanish Latin Romance SVO
-Vietnamese Latin Austroasiatic SVO
-Turkish Latin Turkic SOV
-Russian Cyrillic Slavic SVO*
-Ukrainian Cyrillic Slavic SVO*
-Japanese Kanji Proto-Japonic SOV
-Korean Korean Koreanic SOV
-Hebrew Hebrew Arabic VSO
-Farsi (Persian) Perso-Arabic Indo-Iranian SOV
-Arabic Arabic Arabic VSO
-Table 2: Languages, their scripts, families, and sentence structures (SVO: subject-verb-object, SOV: subject-object-verb, VSO: verb-subject-object, SVO*: SVO dominant but SOV is also possible).
-Appendix B Experimental Setup
+Acknowledgments
+We thank our colleagues at the CoAStaL NLP group Laura Cabello, Rita Ramos, and Israfel Salazar for valuable comments on the final manuscript.
+DE was supported by the European Union’s Horizon 2020 research and innovation program under grant agreement No.
+101135671 (TrustLLM).
+Appendix A Experimental Setup
 We use the mParaRel triplets and templates to query the factual knowledge of the language models.
@@
 Table 3: Total number of examples in mParaRel.
-B.1 Prompt Details
+A.1 Prompt Details
 For mT5, we feed the mParaRel input into the encoder with a sentinel token in the object placeholder.
@@
 Any tokens generated preceding the object, are added to the decoder input since adding these to the encoder does not ensure that the next token predicted will be the object.
-B.2 Computational Resources
+A.2 Computational Resources
 The experiments were run in A100 GPUs. The causal analysis took from 30 minutes to 12 hours depending on the language. The rest of the experiments took 1-2 hours per language.
-Appendix C Causal Tracing
-C.1 XGLM
+Appendix B Causal Tracing
+[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
+Figure 11: Average indirect effect on logit scores, when the subject input has been corrupted and some activations are restored to their uncorrupted values.
+The MHSA and MLP are restored in windows of 12% consecutive layers.
+Top GPT2-XL (only English data).
+B.1 XGLM
 Figure 12: XGLM causal analysis for each language (continues in Figure 13).
@@
 in a window of 4 layers.
-C.2 EuroLLM
+B.2 EuroLLM
 Figure 16: EuroLLM causal analysis for each language (continues in Figure 17).
@@
 in a window of 5 layers.
-C.3 mT5
+B.3 mT5
 Figure 20: mT5 causal analysis for each language (continues in Figure 21).
@@
 in a window of 3 layers.
-Appendix D Attention Knockout
+Appendix C Attention Knockout
 [Uncaptioned image][Uncaptioned image][Uncaptioned image]
@@
 From the encoder sentinel there is no much flow of information so these were not included in the main body.
-Appendix E Extraction Event
+Appendix D Extraction Event
 Figure 28: XGLM extraction rates for each language.
@@
 Figure 34: Number of extraction events split by precedence (or not) of an extraction event in the cross-attn in mT5 for each language.
-Appendix F Patching
+Appendix E Patching
+Figure 35: Aggregated patching results (§6) for XGLM.
+The last token representation contains the function that solves the task.
+This function is formed in two stages: first, the relation to extract is encoded (green), and then the language is composed into the function (purple).
+Finally, the function is applied to the subject, and the predicted object is resolved (blue).
 XGLM EuroLLM mT5
@@
 The total number of examples varies, see total numbers in Table 6.
-F.1 Different Relation, Different Subject
+E.1 Different Relation, Different Subject
 Figure 36: Probability of the patch answer and the context answer when patching at different layers. Models from top to bottom: XGLM, EuroLLM, mT5.
-F.2 Same Relation, Different Subject
+E.2 Same Relation, Different Subject
 Figure 37: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers. Models from top to bottom: XGLM, EuroLLM, mT5.
@@
 Figure 43: Probability of the $\mathcal{L}_{c}(o_{c})$ when patching at different layers in mT5, for examples with $\{\neq\mathcal{L},=r,\neq s\}$. Note that the plots do not share the y-axis.
-F.3 Different Relation, Same Subject
+E.3 Different Relation, Same Subject
 Figure 44: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers for examples with $\{\neq\mathcal{L},\neq r,=s\}$. Models from top to bottom: XGLM, EuroLLM, mT5.